### v1

# On circumcentered direct methods for monotone variational inequality problems

Abstract
The variational inequality problem (VIP) plays a central role in the theory and applications in continuous optimization. In particular, minimization problems and KKT systems can be regard as VIPs. In this work, we present the first methods using circumcenter iterations for solving VIPs. The circumcentered-reflection method (CRM) is a tool based on projections developed with the aim of finding a point in the intersection of finitely many closed convex sets. CRM has gone through enhancements and adaptations over the last few years and was proven to be faster in many settings than competitors such as alternating projections and the Douglas-Rachford method. One of the nice features of CRM is that it is able to deal with approximate projections, which is exactly what we enforced in this article theoretically and numerically. We present both a circumcenter method for the VIP with paramonotone operator and monotone operator, the first employing exact projections and the second approximate ones. Convergence results showing their convergence are established. Numerically, our experiments indicate good performance, including advantages over the well-know extragradient method.
1 Introduction
1.1 Variational inequalities
[…]
$\langle F(x^{*}),x-x^{*}\rangle\geq 0,~{}~{}\forall x\in C.$ (1.1)
The set of solutions of VIP$(F,C)$ will be denoted by SOL$(F,C)$. In particular, if $F=\nabla g$ for a convex function $g:\mathds{R}^{n}\to\mathds{R}$, then VIP$(F,C)$ is equivalent to the problem of minimizing $g$ over $C$.
The variational inequality problem was initially introduced and studied by [36]. Variational inequalities have attracted wide attention, since they cover a large variety of problems arising in different fields, such as economics, structural analysis, optimization, engineering sciences, and operations research, see [2, 20, 22, 34, 35, 44, 36, 51] and the references therein. An excellent survey of methods for variational inequality problems can be found in [30].
1.2 Gradient-type methods for variational inequalities
Here, we are interested in direct methods for solving VIP($F,C$). The basic idea consists of extending the projected gradient method for constrained optimization, i.e., for the problem of minimizing $g(x)$ subject to $x\in C$. This problem is a particular case of VIP($F,C$) with $F=\nabla g$. This procedure is given by the following iterative scheme:
$x^{0}\in C,$ (1.2)
$x^{k+1}=P_{C}(x^{k}-\alpha_{k}\nabla g(x^{k})),$ (1.3)
with $\alpha_{k}>0$ for all $k$. The coefficients $\alpha_{k}$ are called stepsizes and $P_{C}:\mathds{R}^{n}\to C$ is the orthogonal projection onto $C$, i.e., $P_{C}(x)={\rm argmin}_{y\in C}\left\lVert x-y\right\rVert$.
An immediate extension of the method (1.2)–(1.3) to VI$(F,C)$ is the iterative procedure given by
$x^{0}\in C,$ (1.4)
$x^{k+1}=P_{C}(x^{k}-\alpha_{k}F(x^{k})),$ (1.5)
which will be called Gradient-Type (GT) method for variational inequalities.
Convergence results for this method require some monotonicity properties of $F$. We introduce below several possible options.
Definition 1.1.
Consider an operator $F:\mathds{R}^{n}\rightarrow\mathds{R}^{n}$ and a convex set $M\subset\mathds{R}^{n}$. $F$ is said to be
  • (i)
    monotone on $M$ if $\langle F(x)-F(y),x-y\rangle\geq 0$ for all $x,y\in M$;
  • (ii)
    paramonotone on $M$ if it is monotone in $M$, and whenever $\langle x-y,F(x)-F(y)\rangle=0$ with $x,y\in M$, it holds that $F(x)=F(y)$;
  • (iii)
    strictly monotone on $M$ if $\langle x-y,F(x)-F(y)\rangle>0$ for all $x,y\in M$, $x\neq y$;
  • (iv)
    uniformly monotone on $M$ if $\langle x-y,F(x)-F(y)\rangle\geq\varphi(\left\lVert F(x)-F(y)\right\rVert)$ for all $x,y\in M$, where $\varphi:\mathds{R}_{+}\to\mathds{R}$ is an increasing function with $\varphi(0)=0$;
  • (v)
    strongly monotone on $M$ if $\langle x-y,F(x)-F(y)\rangle\geq\omega\left\lVert F(x)-F(y)\right\rVert^{2}$ for all $x,y\in M$ and some $\omega>0$.
Remark: It follows from Definition 1.1 that the following implications hold: (iv) $\Rightarrow$ (iii) $\Rightarrow$ (ii) $\Rightarrow$ (i). The reverse assertions are not true in general.
We mention that all these monotonicity notions can be defined also for point-to-set operators, i.e., for $F:\mathds{R}^{n}\to{\mathcal{P}}(\mathds{R}^{n})$. In this situation, $F$ is usually required to be maximal monotone. In the point-to-point setting (the one considered in this paper), maximal monotonicity reduces to monotonicity and continuity. We make this observation because the proofs of some results that will be invoked later require maximal monotonicity of $F$. Our continuity and monotonicity assumptions ensure that these results hold in our case.
It has been proved in [31] that when $F$ is strongly monotone and Lipschitz continuous, i.e., there exists $L>0$ such that $\left\lVert F(x)-F(y)\right\rVert\leq L\left\lVert x-y\right\rVert$ for all $x,y\in\mathds{R}^{n}$, then the GT method converges to the unique solution of VIP$(F,C)$, provided that $\alpha_{k}\in\left(\epsilon,\frac{2\omega}{L^{2}}\right)$ for all $k$ and for some $\epsilon>0$.
These results were later improved under weaker monotonicity assumptions on the operator $F$. The case of uniformly monotone operators was analyzed in [1] and the case of paramonotone ones in [18]. The first algorithm proposed in this paper is shown to converge when $F$ is paramonotone. Thus, it is worthwhile to look more carefully at this class of operators.
The notion of paramonotonicity was introduced in [21], and many of its properties were established in [23] and [39]. Among them, we mention the following:
  • i)
    If $F$ is the subdifferential of a convex function, then $F$ is paramonotone, see Proposition $2.2$ in [39].
  • ii)
    If $F:\mathds{R}^{n}\to\mathds{R}^{n}$ is monotone and differentiable, and $J_{F}(x)$ denotes the Jacobian matrix of $F$ at $x$, then $F$ is paramonotone if and only if ${\rm Rank}(J_{F}(x)+J_{F}(x)^{t})={\rm Rank}(J_{F}(x))$ for all $x$, see Proposition $4.2$ in [39].
It follows that affine operators of the form $F(x)=Ax+b$ are paramonotone when $A$ is positive semidefinite (not necessarily symmetric), and ${\rm Rank}(A+A^{t})={\rm Rank}(A)$. This situation includes cases of nonsymmetric and singular matrices, where the solution set of VI$(F,C)$ can be a subspace, differently from the case of strictly or strongly monotone operators, for which the solution set is always a singleton, when nonempty. Of course, this can also happen for nonlinear operators, e.g., when $F=\nabla g$ for a convex $g$, in which case the solution set is the set of minimizers of $g$ on $C$, which need not be a singleton. On the other hand, when $A$ is nonzero and skew-symmetric, $F(x)=Ax+b$ is monotone but not paramonotone.
[…]
In order to overcome this difficulty, the following iteration, called extragradient method, was proposed by G.M. Korpelevich in [45]:
$y^{k}=P_{C}\left(x^{k}-\beta F(x^{k})\right),$ (1.6)
$x^{k+1}=P_{C}\left(x^{k}-\beta F(y^{k})\right),$ (1.7)
where $\beta>0$ is a fixed number. Assuming that $F$ is monotone and Lipschitz continuous with constant $L$, that $\beta\in\left(\epsilon,\frac{1}{L}\right)$ for some $\epsilon>0$, and that VIP($F,C$) has solutions, Korpelevich showed that the sequence $\{x^{k}\}$ generated by (1.6)-(1.7) converges to a solution of VIP$(F,C)$.
Our second algorithm, which converges under the sole assumption that $F$ is monotone, is based on this extragradient method.
1.3 The case of $C=\cap_{i=1}^{m}C_{i}$
In principle, these methods are direct, in the sense there is no subproblem to be solved at each iteration, which requires just evaluation of $F$ and computation of the projection $P_{C}$ at certain points. However, the computation of the projection onto a convex set, in general, is not elementary (except for some very special convex sets, such as balls or halfspaces), and hence it is worthwhile to explore situations in which the projection step can be replaced by less expensive procedures. A typical situation is when the convex set $C$ is given as intersections of simpler sets, i.e., $C=\cap_{i=1}^{m}C_{i}$, where each $C_{i}\subset\mathds{R}^{n}$ is closed and convex. For instance, in the very frequent case when $C=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\,(1\leq i\leq m)\}$, with convex $g_{i}:\mathds{R}^{n}\to\mathds{R}$, it is natural to take $C_{i}=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\}$. Usually, $P_{C_{i}}$ will be much easier to compute than $P_{C}$.
A relaxed projection algorithm proposed by He and Yang in [37] for solving VI$(F,C)$ is as follows:
$x^{k+1}=\left(P_{m}^{k}\circ\cdots\circ P_{2}^{k}\circ P_{1}^{k}\right)(x^{k}-\alpha_{k}F(x^{k})).$ (1.8)
where the $P_{i}^{k}$’s are approximations of $P_{C_{i}}$. In the case of exact projections(i.e., $P_{i}^{k}=P_{C_{i}}$) this is a variant of the GT method where the projection onto $C$ is replaced by sequential projections onto the sets $C_{1},\dots C_{m}$. The convergence analysis in [37] requires strong monotonicity of $F$.
Our algorithms also work with easily computable approximate projections onto each set $C_{i}$, but our approach differs from the one in [37] in several directions:
  • i)
    We consider a GT method which converges for paramonotone operators (rather than strongly monotone ones), under a standard coerciveness assumption, and an extragradient-type method which converges for plainly monotone operators
  • ii)
    We use simultaneous (rather than sequential) projections onto the sets $C_{i}$.
  • iii)
    Most importantly, we accelerate the methods through the introduction of a circumcentering procedure, which we explain in the next subsection.
1.4 Circumcentered-Reflection methods
The Convex Feasibility Problem (CFP) consists of finding a point in the intersection of $m$ convex subsets of $\mathds{R}^{n}$.
Two classical methods for CFP are the Sequential and the Simultaneous Projection Methods (SePM and SiPM respectively, which originate in [42] and [24]).
If $C=\cap_{i=1}^{m}C_{i}$, with $C_{i}\subset\mathds{R}^{n}$ closed and convex ($1\leq i\leq m$) and $P_{i}:\mathds{R}^{n}\to C_{i}$ is the orthogonal projection onto $C_{i}$, then the iteration of SePM is of the form
$x^{k+1}=\left(P_{m}\circ\cdots\circ P_{2}\circ P_{1}\right)(x^{k}),$ (1.9)
and the iteration of SiPM is given by
$x^{k+1}=\frac{1}{m}\sum_{i=1}^{m}P_{i}(x^{k})$ (1.10)
The Circumcentered-Reflection Method (CRM), introduced in [14], starts with the case of two convex sets (say $C=A\cap B$) and has the following iteration formula:
$x^{k+1}:=T_{A,B}(x^{k})={\rm circ}(x^{k},R_{A}(x^{k}),R_{B}(R_{A}(x^{k}))),$ (1.11)
where $P_{A}$, $P_{B}$ are the projection operators onto $A,B$, $R_{A}=2P_{A}-{\rm I}$ and $R_{B}=2P_{B}-{\rm I}$ are the standard reflection operators, ${\rm circ}(x^{k},R_{A}(x^{k}),R_{B}(R_{A}(x^{k})))$ is the circumcenter of $x^{k}$, $R_{A}(x^{k})$ and $R_{B}(R_{A}(x^{k}))$, i.e., a point equidistant to these three points and laying in the affine manifold spanned by them. It has been proved in [14] that the sequence generated in this way converges to a point in $C$ if $C$ is nonempty and either $A$ or $B$ is an affine manifold.
However, CRM can be applied to the general CFP with $m$ sets, using Pierra’s product space reformulation (see [49]), which works as follows:
Define ${\textbf{W}},{\textbf{D}}\subset\mathds{R}^{nm}$ as
${\textbf{W}}:=C_{1}\times\cdots\times C_{m}\subset\mathds{R}^{nm},$ (1.12)
and
${\textbf{D}}:=\{{\textbf{y}}\in\mathds{R}^{nm}:\textbf{y}=(x,\ldots,x),x\in\mathds{R}^{n}\}.$ (1.13)
It is immediate that if ${\textbf{y}}=(z,\dots,z)\in\mathds{R}^{nm}$ then
$(P_{\textbf{D}}\circ P_{\textbf{W}})({\textbf{y}})=(\bar{z},\dots,\bar{z})$
[…]
${\rm\mathbf{T}_{\textbf{W},\textbf{D}}}({\textbf{z}})={\rm circ}(\mathbf{R}_{\mathbf{D}}\mathbf{R}_{\mathbf{W}}(\mathbf{z}),\mathbf{R}_{\mathbf{W}}(\mathbf{z}),\mathbf{z}),$ (1.14)
where
$\mathbf{R}_{\mathbf{W}}(\mathbf{z})=\left(R_{C_{1}}(z),\cdots,R_{C_{m}}(z)\right)$ (1.15)
and
$\mathbf{R}_{\mathbf{D}}(\mathbf{z})=\newline \left(2\left(\sum_{i=1}^{m}P_{C_{i}}(z)\right)-z,\ldots,2\left(\sum_{i=1}^{m}P_{C_{i}}(z)\right)-z\right),$ (1.16)
with ${\mathbf{z}}=(z,\dots,z)\in{\mathbf{D}}$. In principle, the operator $\mathbf{T}_{\mathbf{W},\mathbf{D}}$ is defined from $\mathds{R}^{nm}$ to $\mathds{R}^{mn}$. However by Proposition 2.7(i) in [5], we know that if $\mathbf{z}\in\mathbf{D}$, then $\mathbf{T}_{\mathbf{W},\mathbf{D}}(\mathbf{z})\in\mathbf{D}$. Therefore, we conclude that $\mathbf{T}_{\mathbf{W},\mathbf{D}}:\mathbf{D}\rightarrow\mathbf{D}.$
The CRM method has proved to be quite superior to classical projection methods for solving the CRM, achieving superlinear convergence under mild assumptions, while the classical methods are at most linearly convergent; see [5, 4, 10, 11, 12, 13, 6, 7, 9, 8, 27, 46, 48, 47].
1.5 Approximate projections
CRM is superior to most known methods for CFP, but still has a serious drawback: the computation of the projections onto the sets $C_{i}$. A variant of CRM which uses approximate projections in the product space setting was introduced in [3] and denoted as ACRM (Approximate CRM). We begin by defining separating sets for convex sets.
Definition 1.2.
Given a closed convex set $K\subset\mathds{R}^{n}$, a separating operator for $K$ is a point-to-set mapping: $S:\mathds{R}^{n}\to{\mathcal{P}}(\mathds{R}^{n})$ satisfying:
  • i)
    $S(x)$ is closed and convex for all $x\in\mathds{R}^{n}$,
  • ii)
    $K\subset S(x)$ for all $x\in\mathds{R}^{n}$,
  • iii)
    If a sequence $\{z^{k}\}\subset\mathds{R}^{n}$ converges to $z^{*}\in\mathds{R}^{n}$ and $\lim_{k\to\infty}{\rm dist}(z^{k},S(z^{k}))=0$, then $z^{*}\in K$,
  • iv)
    For any convex and compact set $V\subset{\mathbb{R}}^{n}$ there exists $\tau\in(0,1]$ such that
    ${\rm dist}\left(P_{S(x)\cap V}(x),K\right)\leq\tau\,{\rm dist}(x,K),$ (1.17)
where dist(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes the Euclidean distance between a point and a set in $\mathds{R}^{n}$. This notion of separating set was introduced in [3], where only items (i)-(iii) in Definition 1.2 were demanded. In this paper, we additionally require item (iv). We commit a slight abuse of notation, keeping the same definition after adding (iv), which prevents the separating set $S(x)$ from being too close to $x$.
As long as we determine a separating set $S_{i}$ of $C_{i}$ that satisfies the assumptions in Definition 1.2, the exact projection onto $S_{i}$ can be regarded as the approximate projection onto the original set $C_{i}$. Suppose that $S_{i}(x)\subset\mathds{R}^{n}$ is a separating set of $C_{i}$ at point $x\in\mathds{R}^{n}$ for each $i\in\{1,\ldots,m\}$. In order to utilize CRM with approximate projection, we first define the separating set in the product space as follows:
$\mathbf{S}:=S_{1}\times\cdots\times S_{m}\subset\mathds{R}^{nm}.$ (1.18)
Then, the approximate circumcentered-reflection operator $\mathbf{T}_{\mathbf{S},\mathbf{D}}:\mathds{R}^{nm}\rightarrow\mathds{R}^{nm}$ is defined as
$\mathbf{T}_{\mathbf{S},\mathbf{D}}(\mathbf{z})={\rm circ}(\mathbf{R}_{\mathbf{D}}\mathbf{R}_{\mathbf{S}}(\mathbf{z}),\mathbf{R}_{\mathbf{S}}(\mathbf{z}),\mathbf{z}),$ (1.19)
where $circ(\cdot,\cdot,\cdot)$ are the circumcenter of their points as it defined in (1.11), $\mathbf{R}_{\mathbf{S}}$ is defined in the same as in (1.15) but with respect to $\mathbf{S}$ defined in (1.18) and $\mathbf{R}_{\mathbf{D}}$ is defined in (1.16). Therefore, the sequence $\{\mathbf{z}^{k}\}\subset\mathds{R}^{nm}$ generated by ACRM can be written as
$\mathbf{z}^{k+1}=\mathbf{T}_{\mathbf{S}_{k},\mathbf{D}}(\mathbf{z}^{k}),$ (1.20)
where
$\mathbf{S}_{k}=S_{i}^{k}\times\cdots\times S_{m}^{k}\subset\mathds{R}^{nm},$
and $S_{i}^{k}\subset\mathds{R}^{n}$ is any separating set of $C_{i}$ at point $x^{k}$, and $\mathbf{z}^{k}=(x^{k},\ldots,x^{k})$. In view of Proposition 2.7(i) in [5], we conclude that $\mathbf{T}_{\mathbf{S}_{k},\mathbf{D}}:\mathbf{D}:\rightarrow\mathbf{D}$, that is $\{\mathbf{z}^{k}\}\subset\mathbf{D}$ for all $k\geq 1$. Now we go back to $\mathbf{\mathds{R}}^{n}$, and define the approximate circumcentered-reflection operator $T_{\mathbf{S}_{k},\mathbf{D}}:\mathds{R}^{n}\rightarrow\mathds{R}^{n}$ as follows
$T_{\textbf{S}_{k},\textbf{D}}(x)=\bar{x},$ (1.21)
where
$\mathbf{\bar{x}}=\mathbf{T}_{\mathbf{S}_{k},\mathbf{D}}(\mathbf{x}),\qquad\mathbf{\bar{x}}=(\bar{x},\dots,\bar{x})\qquad\mathbf{x}=(x,\dots,x),$ (1.22)
with $\mathbf{\bar{x}},\mathbf{x}\in\mathds{R}^{nm}$ and $\bar{x},x\in\mathds{R}^{n}$. Consequently, the sequence $\{x^{k}\}\subset\mathds{R}^{n}$ generated by ACRM is given by
$x^{k+1}=T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}).$ (1.23)
In practice, the sets $S^{k}_{i}$ will be halfspaces, so that the projections onto the sets $S^{k}_{i}$ are easily computable. The same holds for the operators $R_{\textbf{W}},R_{{\textbf{S}}_{k}}$, and consequently for the operator $T_{{\textbf{S}}_{k},{\textbf{D}}}$. This allows us to obtain a closed formula for $x^{k+1}$ in terms of $x^{k}$ in (1.21). This is the case, for instance, of Algorithm PACA (see [11]), which, using suitably perturbed separating sets, generates a sequence that is finitely convergent to a solution of CFP, under a Slater condition.
2 Preliminaries
We now present some previously published results which will be used in our analysis.
Proposition 2.1.
Assume that $K\subset\mathds{R}^{n}$ is a closed convex set. Then,
  • i)
    $\langle x-P_{K}(x),y-P_{K}(x)\rangle\leq 0,~{}\forall x\in\mathds{R}^{n}~{}and~{}\forall y\in K.$
  • ii)
    $\left\lVert x-y\right\rVert^{2}\geq\left\lVert x-P_{K}(x)\right\rVert^{2}+\left\lVert y-P_{K}(x)\right\rVert^{2}$.
Proof.
See Lemma 2 in [17]. ∎
Proposition 2.2.
If $H=\{y\in\mathds{R}^{n}\mid a^{\top}y\leq\alpha\}$, with $a\in\mathds{R}^{n}$, $\alpha\in R$, then the orthogonal projection onto $H$ is given by
$P_{H}(x)=x-\frac{\max\{0,\alpha-a^{\top}x\}}{\|a\|^{2}}a.$
Proof.
See Proposition 2.2 in [11]. ∎
Proposition 2.3.
Suppose $C\subset\mathds{R}^{n}$ is a closed convex set, $U\subset\mathds{R}^{n}$ is an affine subspace. Let $S\subset\mathds{R}^{n}$ be a separating set of $C$ which satisfies $(i)-(iii)$ in Definition 1.2. Define $H(x)\subset\mathds{R}^{n}$ as:
$H(x)=\begin{cases}K,&{\rm if}\,\,\,x\in C,\\ \{z\in\mathds{R}^{n}:\langle z-P_{S}(x),x-P_{S}(x)\rangle\leq 0\},&{\rm otherwise}.\end{cases}$ (2.1)
Then, for all $x\in U$, we have
$T_{S,U}(x)=P_{H(x)\cap U}(x).$ (2.2)
Proof.
See Proposition 3.3 in [3]. ∎
Next, we present an explicit instance of the approximate circumcenter operator $T_{\mathbf{S},\mathds{D}}$, defined in (1.21). First, we define a separating set $S_{i}(y)\subset\mathds{R}^{n}$ associated to each set $C_{i}$, assumed to be of the form:
$C_{i}=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\},$ (2.3)
with a convex $g_{i}:\mathds{R}^{n}\to\mathds{R}$. $S_{i}$ is defined as follows:
$S_{i}(y)=\begin{cases}C_{i}&{\rm if}\,\,y\in C_{i},\\ \{z\in\mathds{R}^{n}:(u_{i})^{\top}(z-y)+g_{i}(x)\leq 0\},&{\rm otherwise},\end{cases}$ (2.4)
where $u_{i}$ is a subgradient of $g_{i}$ at $x^{k}$. Given ${\mathbf{y}}=(y^{1},\dots,y^{m})\in\mathds{R}^{nm}$, we define ${\mathbf{S}}({\mathbf{y}})\subset\mathds{R}^{nm}$ as
${\mathbf{S}}({\mathbf{y}})=S_{1}(y^{1})\times\dots\times S_{m}(y^{m}).$ (2.5)
We have the following result related to these separating sets.
Proposition 2.4.
Suppose that $C_{i}\subset\mathds{R}^{n}$ and $S_{i}(y)\subset\mathds{R}^{n}$ are defined as in (2.3) and (2.4). Then $C_{i}\subset S_{i}(y)$ for any $y\in\mathds{R}^{n}$ and $S_{i}(y)$ is a separating set for $C_{i}$ at the point $y$ that satisfies conditions (i)-(iii) in Definition 1.2. Also, ${\mathbf{S}}({\mathbf{y}})$ is a separating set for $C_{1}\times\dots\times C_{m}\subset\mathds{R}^{nm}$ at ${\mathbf{y}}$.
Proof.
See Proposition 2.8 in [5]. ∎
Remark 2.5.
Take $C_{i}$ as in (2.3), $T_{\mathbf{S}_{k},\mathbf{D}}$ as in (1.21) and $\mathbf{S}_{k}$ as above. By Proposition 2.4 and 2.3, we immediately have $T_{\mathbf{S}_{k},\mathbf{D}}(x)=P_{\mathbf{S}_{k}\cap\mathbf{D}}(x)=P_{\mathbf{H}_{k}\cap\mathbf{D}}(x)$, where $\mathbf{H}_{k}$ is defined in the same way as $H(x)$ in (2.1) with respect to the separating set $\mathbf{S}_{k}$.
Proposition 2.6.
Assume that $K\subset\mathds{R}^{n}$ is a convex and closed set and $U\subset\mathds{R}^{n}$ is an affine subspace. Then
$\left\lVert T_{K,D}(x)-y\right\rVert\leq\left\lVert T_{MAP}(x)-y\right\rVert,$ (2.6)
for all $x\in U$, for all $y\in K\cap U$ and $T_{K,D}$ is the circumcenter operator as defined in (1.11) and $T_{MAP}=P_{U}\circ P_{K}$ is the alternating projection operator with respect to $K$ and $U$.
Proof.
See Proposition 2.7 (ii) in [5]
Proposition 2.7.
If $F:\mathds{R}^{n}\to\mathds{R}^{n}$ is continuous then SOL$(T,C)$ is closed.
Proof.
Follows the definition of VIP$(F,C)$ and the continuity of $F$. ∎
Proposition 2.8.
Let $F:\mathds{R}^{n}\to\mathds{R}^{n}$ be a continuous and paramonotone operator. Let $\{z^{k}\}$ be a bounded sequence such that all cluster points of $\{z^{k}\}$ belong to $C$. For each $x\in{\rm SOL}(F,C)$ define $\gamma_{k}:=\langle F(x^{k}),z^{k}-x\rangle$. If for some $\bar{x}\in{\rm SOL}(F,C)$ there exists a subsequence $\{\gamma_{k_{j}}\}$ of $\{\gamma_{k}\}$ such that $\lim_{j\to\infty}\gamma_{k_{j}}\leq 0$, then there exists a cluster point of $\{z^{k_{j}}\}$ belonging to ${\rm SOL}(F,C)$.
Proof.
See Lemma 6 in [17]. ∎
The previous two propositions were stated and proved in [17] for point-to-set operators. We have rephrased their statements to fit our point-to-point setting.
Proposition 2.9.
Take $\{\varepsilon_{k}\}$, $\{v_{k}\}$subset $\mathds{R}_{+}$ and $\mu\in[0,1)$. If the inequalities
$\varepsilon_{k+1}\leq\mu\varepsilon_{k}+v_{k},~{}~{}k\in\mathds{N}$ (2.7)
hold and $\lim_{k\rightarrow\infty}v_{k}=0$, then $\lim_{k\rightarrow}\varepsilon_{k}=0$.
Proof.
See Lemma 2 in [33]. ∎
Proposition 2.10.
Take $K\subset\mathds{R}^{n}$ closed and $z\notin K$. Let $\{z^{k}\}\subseteq\mathds{R}^{n}$ be such that $\lim_{k\rightarrow\infty}\|z^{k+1}-z^{k}\|=0$ and both $z$ and some point in $K$ are cluster points of $\{z^{k}\}$. Then there exist $\zeta>0$ and a subsequence $\{z^{k_{j}}\}$ of $\{z^{k}\}$ such that
${\rm dist}(z^{k_{j}+1},K)>{\rm dist}(z^{k_{j}},K)$ (2.8)
and
${\rm dist}(z^{k_{j}},K)>\zeta.$ (2.9)
Proof.
See Lemma 3 in [17]. ∎
Definition 2.11.
We say $w\in\mathds{R}^{n}$ is a Slater point for ${\rm SOL}(F,C)$ if
$g_{i}(w)<0,$ (2.10)
for all $i=1,\ldots,m$.
Definition 2.12.
Let $S\subset\mathds{R}^{n}$. A sequence $\{x^{k}\}\subset\mathds{R}^{n}$ is called quasi-Fejér convergent to $S$ if for all $x\in S$ there exist $k_{0}\geq 0$ and a sequence $\{\delta_{k}\}\subset\mathds{R}^{n}_{+}$ such that $\sum_{k=0}^{\infty}\delta_{k}<\infty$ and
$\|x^{k+1}-x\|^{2}\leq\|x^{k}-x\|^{2}+\delta_{k},$ (2.11)
for all $k\geq k_{0}$.
Proposition 2.13.
Let $F:\mathds{R}^{n}\rightarrow\mathds{R}^{n}$ be monotone and continuous and $C$ a closed and convex set. Then ${\rm SOL}(F,C)$ is closed and convex.
Proof.
See Lemma 2.4(ii) in [16]. ∎
Proposition 2.14.
Consider the variational inequality problem $\rm{VIP}(F,C)$. If $F:\mathds{R}^{n}\rightarrow\mathds{R}^{n}$ is continuous and monotone, then
${\rm SOL}(F,C)=\{x\in C:\langle F(y),y-x\rangle\geq 0,~{}\forall y\in C\}.$ (2.12)
Proof.
See Lemma 3 in [18]. ∎
Proposition 2.15.
If $\{x^{k}\}$ is a quasi-Fejér sequence with respect to $S$ then:
  • (i)
    $\{x^{k}\}$ is bounded;
  • (ii)
    $\{\|x^{k}-x\|\}$ converges for all $x\in S$;
  • (iii)
    if all cluster point of $\{x^{k}\}$ belong to $S$, then the sequence $\{x^{k}\}$ converges to a point in $S$.
Proof.
See Proposition 1 in [18]. ∎
Proposition 2.16.
If a sequence $\{x^{k}\}$ is quasi-Fejér convergent to a closed and convex set $S\subset\mathds{R}^{n}$ then
  • (i)
    the sequence $\{dist(x^{k},S)\}$ is convergent;
  • (ii)
    the sequence $\{P_{S}(x^{k})\}$ is convergent.
Proof.
See Lemma 2 in [18]. ∎
Proposition 2.17.
Let $\{u^{k}\}\subset\mathds{R}^{n}$ be a sequence which converges to $\bar{u}$. Take nonnegative real number $\zeta_{k,j}(k\geq 0,0\leq j\leq k)$ such that $\lim_{k\rightarrow\infty}\zeta_{k,j}=1$ for all $k$. Define
$w^{k}:=\sum_{j=0}^{k}\zeta_{k,j}u^{j}.$ (2.13)
Then, $\{w^{k}\}$ also converges to $\bar{u}$.
Proof.
Elementary. ∎
Lemma 2.18.
Let $g_{i}:\mathds{R}^{n}\rightarrow\mathds{R}$ is convex function for all $i\in\{1,\ldots,m\}$ and $C_{i}:=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\}$. Define $C:=\cap_{i=1}^{m}C_{i}$. Assume that there exists a Slater point $w\in C$ such that $g(w)<0$ where $g$ is defined in (4.5). Then, for all $y$ such that $g(y)>0$, we have
$\operatorname*{dist}(y,C)\leq\frac{\|y-w\|}{g(y-g(w))}g(y).$ (2.14)
Proof.
Take $w_{\lambda}:=\lambda w+(1-\lambda)y$ with $\lambda=\frac{g(y)}{g(y)-g(w)}$. Note that $\lambda\in(0,1)$. Then
$g(w_{\lambda})=g(\lambda w+(1-\lambda)y)\leq\lambda g(w)+(1-\lambda)g(y)=g(y)-\lambda(g(y)-g(w))=0,$ (2.15)
where the first inequality holds because $g$, as defined in (4.5), is a convex function. Thus, $w_{\lambda}\in C$ and
$dist(y,C)\leq\|y-w_{\lambda}\|=\|y-(\lambda w+(1-\lambda)y)\|=\lambda\|y-w\|=\frac{g(y)}{g(y)-g(w)}\|y-w\|.$ (2.16)
3 A circumcenter method for the variational inequality problem with paramonotone operator
In this paper, we focus on the variational inequality problem VIP$(F,C)$ with a feasible set of the form $C:=\cap_{i=1}^{m}C_{i}$, where each $C_{i}\subset\mathds{R}^{n}$ is closed and convex, i.e., we deal with the problem:
${\rm find}~{}x^{*}\in C\\ \,\,\,{\rm such\,\,\,that}\,\,\,\langle F(x^{*}),x-x^{*}\rangle\geq 0\,\,\,\forall x\in C.$ (3.1)
We propose two different algorithms, depending on the monotonicity and continuity properties of $F$. Therefore, we list next several assumptions, some of which will be needed for the convergence analysis of each algorithm.
Assumption 3.1.
The solution set of Problem (3.1), i.e., SOL$(F,C)$, is nonempty.
Assumption 3.2.
The mapping $F$ is Lipschitz continuous on $C$ with some constant $L>0$, i.e.,
$\left\lVert F(x)-F(y)\right\rVert\leq L\left\lVert x-y\right\rVert,\,\,\,\forall x,y\in C.$ (3.2)
Assumption 3.3.
The mapping $F$ is monotone on $C$.
Assumption 3.4.
The mapping $F$ is paramonotone on $C$.
Assumption 3.5.
The mapping $F$ satisfies the following coerciveness condition: there exist $\hat{z}\in C$ and a bounded set $W\subset\mathds{R}^{n}$ such that $\langle F(x),x-\hat{z}\rangle\geq 0$ for all $x\notin W$.
Some comments are in order regarding Assumption 3.5. Coerciveness assumptions have been required in previous convergence analysis of direct methods for variational inequalities, for instance in [33], where a condition stronger than Assumption 3.5 is demanded, and in [17], where Assumption 3.5 is used. Our assumption holds immediately when $F$ is monotone and has a zero $\bar{z}$, in which case we take $\hat{z}=\bar{z}$ and $D=\emptyset$, and also when $F$ is uniformly monotone with $\lim_{t\to\infty}\psi(t)/t=\infty$ (for more details, see [17], where other monotone operators satisfying this assumption are discussed).
In this section, we consider a one-step circumcenter method, based on the GT method, for solving VI$(F,C)$ with a paramonotone operator $F$ and a feasible set of the form $C=\cap_{i=1}^{m}$.
The starting point for our method will be Algorithms 1 and 2 in [17], which are versions of the TG method for point-to-set paramonotone operators. We will add the circumcenter approach introduced in [14] and [15].
3.1 The algorithm
We present next the one-step circumcenter method for solving VIP$(F,C)$, with paramonotone $F$ and feasible set $C$ of the form $C=\cap_{i=1}^{m}C_{i}$.
Algorithm 1 Approximate Circumcenter method for VIP$(F,C)$ with paramonotone $F$
Initialization step
: Choose $x^{0}\in\mathds{R}^{n}$.
Iterative step
: Given $x^{k}$, if $F(x^{k})=0$ then stop. Otherwise, define $\eta_{k}:=\max\{1,\left\lVert F(x^{k})\right\rVert\}$. Choose a separating operator $S^{i}:\mathds{R}^{n}\to{\mathcal{P}}(\mathds{R}^{n})$ for the set $C_{i}$, define $S^{i}_{k}:=S^{i}(x^{k})$ and ${\mathbf{S}}_{k}:=S^{1}_{k}\times\dots\times S^{m}$.Then, define the next iterate as:
$x^{k+1}=T_{\textbf{S}_{k},\textbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right),$ (3.3)
where $T_{\textbf{S}_{k},\textbf{D}}$ is the approximate circumcentered-reflection operator defined in (1.23), and the stepsize sequence $\{\beta_{k}\}$ satisfies that:
$\sum_{k=0}^{\infty}\beta_{k}=\infty,$ (3.4)
and
$\sum_{k=0}^{\infty}\beta_{k}^{2}<\infty,$ (3.5)
We remark that when $F$ is the gradient of a smooth convex function $f:\mathds{R}^{n}\to\mathds{R}$ (in which case, as mentioned above, $F$ is indeed paramonotone), Algorithm 1 can be seen as an upgraded Projected Gradient method, with the acceleration effect resulting from the circumcenter approach.
3.2 Convergence Analysis
We need some preliminary propositions for our convergence theorem.
Proposition 3.6.
Suppose that $C=\cap_{i=1}^{m}C_{i}$. Let $V\subset\mathds{R}^{n}$ be a nonempty compact and convex set. Define ${\mathbf{V}}=:V\times\cdots\times V\subset\mathds{R}^{nm}$. Then, there exist $\bar{\tau}\in[0,1)$ such that
${\rm dist}(T_{{\mathbf{S}}_{k}\cap{\mathbf{X}},{\mathbf{D}}}(x),C)\leq\bar{\tau}\,{\rm dist}(x,C),$ (3.6)
for all $x\in V\setminus C$, where $T_{{\mathbf{S}}_{k}\cap{\mathbf{X}},{\mathbf{D}}}$ is as (1.21) and ${\mathbf{S}_{k}}=S_{k}^{1}\times\dots\times S^{m}_{k}$, where $S^{i}_{k}=S_{i}(x^{k})$ and $S^{i}:\mathds{R}^{n}\to{\mathcal{P}}(\mathds{R}^{n})$ is a separating set for $C_{i}$ in the sense of Definition 1.2.
Proof.
First, we have
${\rm dist}(T_{\mathbf{S}_{k}\cap\mathbf{V},\mathbf{D}}(x),C)\leq{\rm dist}((P_{\textbf{D}}\circ P_{\mathbf{S}_{k}\cap\mathbf{V}})(x),C),$ (3.7)
according to Proposition 2.6 with $K=\mathbf{S}_{k}\cap\mathbf{V}$ and $U=\mathbf{D}$. Now we invoke Definition 1.2(iv) for establishing that there exists $\tau_{i}\in(0,1]$ such that
${\rm dist}(P_{S_{i}\cap V}(x),C)\leq\tau_{i}\,{\rm dist(x,C)}$ (3.8)
Define $\bar{\tau}=\max{\tau_{1},\dots,\tau_{m}}$. Using the definition of $\mathbf{D}$ in (1.13), we have
$\begin{split}{\rm dist}((P_{\textbf{D}}\circ P_{\mathbf{S}_{k}\cap\mathbf{V}})(x),C)&={\rm dist}\left(\frac{1}{m}\sum_{i=1}^{m}P_{S_{i}\cap V}(x),C\right)\\ &\leq\frac{1}{m}\sum_{i=1}^{m}{\rm dist}(P_{S_{i}\cap V}(x),C)\\ &\leq\frac{1}{m}\sum_{i=1}^{m}\tau_{i}\,{\rm dist(x,C)}\\ &\leq\bar{\tau}\,{\rm dist}(x,C),\end{split}$ (3.9)
for all $x\in V\setminus C$. The first inequality in (3.9) holds by convexity the distance function, the second one holds by (3.8) and the third one by the definition of $\bar{\tau}$. Using (3.7) and (3.9), we complete the proof. ∎
Lemma 3.7.
Let $C=\cap_{i=1}^{m}C_{i}$. Consider the sequence $\{x^{k}\}$ defined by Algorithm 1. If $x^{k+1}=x^{k}$ for some $k$, then $x^{k}\in{\rm SOL(F,C)}$.
Proof.
We suppose that $x^{k+1}=x^{k}$, so in the product space $\mathds{R}^{nm}$ we have $\mathbf{x}^{k+1}=\mathbf{x}^{k}$. Since $\mathbf{x}^{k+1}=T_{\textbf{S}_{k},\textbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)$, we get from Proposition 2.3 that $\mathbf{x}^{k}\in\mathbf{H}^{k}\subset\mathbf{S}^{k}\subset\mathds{R}^{nm}$. Therefore, $x^{k}\in S_{i}^{k}$ for all $i\in\{1,\ldots,m\}$. Given $j\in\{1,\ldots,m\}$, we have two cases according to the definition of $S_{i}^{k}$ in (2.4). If $S_{i}^{k}=C_{i}$, then we have $x^{k+1}=x^{k}\in C_{j}$. Otherwise, we have $g_{j}(x^{k})\leq g_{j}(x^{k})+\langle v^{jk},x^{k+1}-x^{k}\rangle\leq 0$, i.e., $x^{k}\in C_{j}$. By the above argument above, we conclude that for any $i\in\{1,\ldots,m\}$, we have $x^{k}\in C_{i}$. Therefore, $x^{k}\in C=\cap_{i=1}^{m}C_{i}$. Moreover, since $x^{k+1}$ is given by (3.3), using Proposition 2.1 (i) in the product space, with $x=\mathbf{x}^{k}-\frac{\beta_{k}}{\eta_{k}}T_{\mathbf{S}_{k},\mathbf{D}}(\mathbf{x}^{k})$, the approximate circumcenter operator coincides with the projection operator, in view of Proposition 2.3, i.e., $T_{\mathbf{S}_{k},\mathbf{D}}(x)=P_{\mathbf{S}_{k}\cap\mathbf{D}}(x)$, and $K=\mathbf{S}_{k}$. Hence we obtain (first in the product space)
$\left\langle\mathbf{x}^{k+1}-(\mathbf{x}^{k}-\frac{\beta_{k}}{\eta_{k}}F(\mathbf{x}^{k})),\mathbf{z}-\mathbf{x}^{k+1}\right\rangle\geq 0~{}~{}\forall\mathbf{z}\in\mathbf{S}_{k},$ (3.10)
and then, in ${\mathbf{R}}^{n}$,
$m\left\langle x^{k+1}-(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})),z-x^{k+1}\right\rangle\geq 0~{}~{}\forall z\in S_{i}^{k},$ (3.11)
for all $i\in\{1,\ldots,m\}$. Taking $x^{k+1}=x^{k}$ in (3.11) and taking into account the facts that $\beta_{k}>0,\eta_{k}\geq 1$ for all $k$ and $z\in C=\cap_{i=1}^{m}C_{i}\subset S_{i}^{k}$ in Proposition 2.4, we get $\langle F(x^{k}),z-x^{k}\rangle\geq 0$ for all $z\in C$. Then, we conclude that $x^{k}\in{\rm SOL(F,C)}$. ∎
The following technical lemma will be used for establishing boundedness of $\{x^{k}\}$.
Lemma 3.8.
Take $\hat{z}\in C:=\cap_{i=1}^{m}C_{i}$ and $W\subset{\mathbb{R}}^{n}$ as in Assumption 3.5. Let $\{x^{k}\}$ be a sequence generated by Algorithm 1. Choose $\lambda>0$ such that $\left\lVert x^{0}-\hat{z}\right\rVert\leq\lambda$ and $W\subset B(\hat{z},\lambda)$. Then
  • (i)
    if $x^{k}\in W$ then $\left\lVert x^{k+1}-\hat{z}\right\rVert^{2}\leq\lambda^{2}+\beta_{k}^{2}+2\beta_{k}\lambda$,
  • (ii)
    if $x^{k}\notin W$ then $\left\lVert x^{k+1}-\hat{z}\right\rVert^{2}\leq\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\beta_{k}^{2}$.
Proof.
Since $\hat{z}\in C$, we get from (3.42) that $\hat{z}=T_{\mathbf{S}_{k},\mathbf{D}}(\hat{z})$. Then, in view of (3.3) and Proposition 2.1(i), we obtain
$\begin{split}\left\lVert x^{k+1}-\hat{z}\right\rVert^{2}&=\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k}))-\hat{z}\right\rVert^{2}=\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(\hat{z})\right\rVert^{2}\\ &\leq\left\lVert\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)-\hat{z}\right\rVert^{2}\\ &=\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\frac{\beta_{k}^{2}}{\eta_{k}^{2}}\left\lVert F(x^{k})\right\rVert^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(x^{k}),x^{k}-\hat{z}\rangle\\ &\leq\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(x^{k}),x^{k}-\hat{z}\rangle.\end{split}$ (3.12)
Now,
  • (i)
    if $x^{k}\in W$, applying Cauchy-Schwartz inequality in (3.12), the definition of $\eta_{k}$ and the fact that $W\subset B(\hat{z},\lambda)$, we conclude that
    $\left\lVert x^{k+1}-\hat{z}\right\rVert^{2}\leq\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\beta_{k}^{2}+2\frac{\beta_{k}}{\eta_{k}}\left\lVert F(x^{k})\right\rVert\left\lVert x^{k}-\hat{z}\right\rVert\leq\lambda^{2}+\beta_{k}^{2}+2\beta_{k}\lambda.$ (3.13)
  • (ii)
    if $x^{k}\notin W$, it follows from Assumption 3.5 that $\langle F(x^{k}),x^{k}-\hat{z}\rangle\geq 0$, and we get from (3.13) that
    $\left\lVert x^{k+1}-\hat{z}\right\rVert^{2}\leq\left\lVert x^{k}-\hat{z}\right\rVert^{2}+\beta_{k}^{2},$ (3.14)
    using the fact that $\frac{\beta_{k}}{\eta_{k}}>0$.
Lemma 3.9.
Let $\{x^{k}\}$, $\{F(x^{k})\}$ be the sequences generated by Algorithm 1. Then,
  • (i)
    $\{x^{k}\}$ and $\{F(x^{k})\}$ are bounded,
  • (ii)
    $\lim_{k\rightarrow\infty}\operatorname*{dist}(x^{k},C)=0$,
  • (iii)
    $\lim_{k\rightarrow\infty}\|x^{k+1}-x^{k}\|=0$,
  • (iv)
    All cluster points of $\{x^{k}\}$ belong to $C$.
Proof.
  • (i)
    Take $\hat{z}$ and $W$ as in Assumption 3.5, $\lambda>0$ such that $\left\lVert x^{0}-\hat{z}\right\rVert\leq\lambda$ and $W\subset B(\hat{z},\lambda)$, and $\bar{\beta}>0$ such that $\beta_{k}\leq\bar{\beta}$ for all $k$ (such a $\bar{\beta}$ exists because $\sum_{k=0}^{\infty}\beta_{k}^{2}<\infty$). Let $\sigma=\sum_{j=0}^{\infty}\beta_{j}^{2}$. Define $\bar{\lambda}:=(\lambda^{2}+\bar{\beta}\lambda+\sigma)^{\frac{1}{2}}$. We claim that
    $\{x^{k}\}\subset B(\hat{z},\bar{\lambda}).$ (3.15)
    We proceed to prove the claim. If $x^{k}\in B(\hat{z},\lambda)$, we have $x^{k}\in B(\hat{z},\bar{\lambda})$, since $\bar{\lambda}>\lambda$. Otherwise, i.e., if $x^{k}\notin B(\hat{z},\lambda)$, let $\ell(k)=\max\{\ell<k:x^{\ell}\in B(\hat{z},\bar{\lambda})\}$. Observe that $\ell(k)$ is well-defined because $\|x^{0}-\hat{z}\|\leq\lambda$, so that $x^{0}\in B(\hat{z},\bar{\lambda})$. By Lemma 3.8(i),
    $\left\lVert x^{\ell(k)+1}-\hat{z}\right\rVert^{2}\leq\lambda^{2}+\beta_{\ell(k)}^{2}+2\beta_{\ell(k)}\lambda\leq\lambda^{2}+\bar{\beta}\lambda+\beta_{\ell(k)}^{2}.$ (3.16)
    Iterating the inequality in Lemma 3.8(ii), since $x^{j}\notin W$ for $j$ between $\ell(k)+1$ and $k-1$, we obtain
    $\left\lVert x^{k}-\hat{z}\right\rVert^{2}\leq\left\lVert x^{\ell(k)+1}-\hat{z}\right\rVert^{2}+\sum_{j=\ell(k)+1}^{k-1}\beta_{k}^{2}..$ (3.17)
    Combining (3.16) and (3.17)
    $\begin{split}\left\lVert x^{k}-\hat{z}\right\rVert^{2}&\leq\lambda^{2}+2\bar{\beta}\lambda+\sum_{j=\ell(k)}^{k-1}\beta_{j}^{2}\\ &\leq\lambda^{2}+2\bar{\beta}\lambda+\sum_{j=0}^{\infty}\beta_{j}^{2}\\ &=\lambda^{2}+2\bar{\beta}\lambda+\sigma\\ &=\bar{\lambda}^{2}.\end{split}$ (3.18)
    Thus, $x^{k}\in B(\hat{z},\bar{\lambda})$, proving the claim, and hence $\{x^{k}\}$ is bounded. For $\{F(x^{k})\}$, use the boundedness of $\{x^{k}\}$ and Proposition 2.7(iii).
  • (ii)
    For all $k\geq 1$ we have that
    $\left\lVert x^{k+1}-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert=\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert\leq\frac{\beta_{k}}{\eta_{k}}\left\lVert F(x^{k})\right\rVert\leq\beta_{k},$ (3.19)
    using (3.3) and nonexpansiveness of $T_{\mathbf{S}_{k},\mathbf{D}}$ in the first inequality, and the fact that $\eta_{k}\geq\|F(x^{k})\|$ for all $k$ in the second one.
    Since we already proved that $x^{k}\in B(\hat{z},\bar{\lambda})$ for all $k\geq 1$, we conclude that
    $T_{\mathbf{S}_{k}\cap\mathbf{X},\mathbf{D}}(x)=T_{\mathbf{S}_{k},\mathbf{D}}(x)$ (3.20)
    for any $x\in B(\hat{z},\lambda)\setminus C$. Now we invoke Proposition 3.6 with $V=B(\hat{z},\bar{\lambda})$ and conclude that there exists $\tilde{\mu}\in[0,1)$ such that
    ${\rm dist}(T_{\mathbf{S}_{k},\mathbf{D}}(x),C)\leq\tilde{\mu}\,{\rm dist}(x,C),$ (3.21)
    for all $x\in B(\hat{z},\bar{\lambda})\setminus C$.
    By (3.15), $\{x^{k}\}\subset B(\hat{z},\bar{\lambda})$, so that it follows from (3.21) that
    ${\rm dist}(T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}),C)\leq\tilde{\mu}\,{\rm dist}(x^{k},C),$ (3.22)
    for all $k$ such that $x^{k}\notin C=\cap_{i=1}^{m}C_{i}$. If $x^{k}\in C$, (3.22) holds trivially. Otherwise, observe that
    $\begin{split}{\rm dist}(x^{k+1},C)&\leq\left\lVert x^{k+1}-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert+{\rm dist}(T_{\mathbf{S}_{k},\mathbf{D}}(x^{k}),C)\\ &\leq\beta_{k}+\tilde{\mu}\,{\rm dist}(x^{k},C),\end{split}$ (3.23)
    using (3.23) and (3.21) in the second inequality. Therefore, using Proposition 2.9 with $v_{k}=\beta_{k}$, $\mu=\tilde{\mu}$ and $\varepsilon_{k}={\rm dist}(x^{k},C)$, we obtain
    $\lim_{k\to\infty}{\rm dist}(x^{k},C)=0,$ (3.24)
    establishing (ii).
  • (iii)
    Using (3.19), we get
    $\left\lVert x^{k+1}-x^{k}\right\rVert\leq\left\lVert x^{k+1}-T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})\right\rVert+\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}(x^{k})-x^{k}\right\rVert\leq\beta_{k}+{\rm dist}(x^{k},C).$ (3.25)
    Since $\lim_{k\to\infty}\beta_{k}=0$ by (3.4), it follows from (ii) and (3.25) that $\lim_{k\to\infty}\left\lVert x^{k+1}-x^{k}\right\rVert$.
  • (iv)
    Follows from (ii).
Paramonotonicity of $F$ is used for the first time in this section in the following theorem.
Theorem 3.10.
If $F$ is paramonotone and SOL$(F,C)\neq\emptyset$, then all cluster points of any sequence $\{x^{k}\}$ generated by Algorithm 1 solve VIP$(F,C)$.
Proof.
Let $\{x^{k}\}$, $\{F(x^{k})\}$ be sequences generated by Algorithm 1. Define $\gamma_{k}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ as
$\gamma_{k}(x):=\langle F(x^{k}),x^{k}-x\rangle.$ (3.26)
Take any $\bar{x}\in{\rm SOL}(F,C)$. From Proposition 2.1(i) and the definition of $\eta_{k}$, we get that
$\begin{split}\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}&=\left\lVert T_{\mathbf{S}_{k},\mathbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(\bar{x})\right\rVert^{2}\\ &\leq\left\lVert\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}\right)-\bar{x}\right\rVert^{2}\\ &=\left\lVert x^{k}-\bar{x}\right\rVert^{2}+\frac{\beta_{k}^{2}}{\eta_{k}^{2}}\left\lVert F(x^{k})\right\rVert^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(x^{k}),x^{k}-\bar{x}\rangle\\ &\leq\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\beta_{k}\left(2\frac{\gamma_{k}(\bar{x})}{\eta_{k}}-\beta_{k}\right).\end{split}$ (3.27)
First we prove that $\{x^{k}\}$ has cluster points in SOL$(F,C)$ Since $\{(x^{k}\},\{F(x^{k}))\}$ are bounded by Lemma 3.9, in view of Proposition 2.8 it suffices to prove that $\{\gamma_{k}(\bar{x})\}$ has a nonpositive cluster point. Assume that this is not true. Our assumption implies that $\{\gamma(\bar{x})\}$ must be bounded away from zero for large $k$, i.e., there exist $\bar{k}$ and $\rho>0$ such that $\gamma_{k}(\bar{x})\geq\rho$ for all $k\geq\bar{k}$. Since $\{F(x^{k})\}$ is bounded, there exists $\theta>1$ such that $\left\lVert F(x^{k})\right\rVert\leq\theta$ for all $k$. Therefore,
$\eta_{k}=\max\{1,\left\lVert u^{k}\right\rVert\}\leq\max\{1,\theta\}=\theta,$ (3.28)
for all $k$. Thus, we can find $\bar{\rho}>0$ such that
$\frac{\gamma_{k}(\bar{x})}{\eta_{k}}\geq\frac{\gamma_{k}(\bar{x})}{\theta}>\bar{\rho},$
and hence, in view of (3.27), we obtain
$\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}\leq\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\beta_{k}(2\rho-\beta_{k}),$ (3.29)
Since $\lim_{k\to\infty}\beta_{k}=0$ by (3.4), there exists $k^{\prime}\geq\bar{k}$ such that $\beta_{k}\leq\bar{\rho}$ for all $k\geq k^{\prime}$. So, we get from (3.29), that, for all $k\geq k^{\prime}$,
$\bar{\rho}\beta_{k}\leq\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}.$ (3.30)
Summing (3.30) with $k$ between $k^{\prime}$ and $q$, we obtain:
$\begin{split}\bar{\rho}\sum_{k=k^{\prime}}^{q}\beta_{k}&\leq\sum_{k=k^{\prime}}^{q}\left(\left\lVert x^{k}-\bar{x}\right\rVert^{2}-\left\lVert x^{k+1}-\bar{x}\right\rVert^{2}\right)\\ &\leq\left\lVert x^{k^{\prime}}-\bar{x}\right\rVert^{2}-\left\lVert x^{q+1}-\bar{x}\right\rVert^{2}\\ &\leq\left\lVert x^{k^{\prime}}-\bar{x}\right\rVert^{2}.\end{split}$ (3.31)
Taking limits in (3.31) with $q\to\infty$, we contradict assumption (3.4). Thus, there exists a cluster point of $\{x^{k}\}$ belonging to SOL$(F,C)$.
Finally, we prove that all cluster points of $\{x^{k}\}$ belong to SOL$(F,C)$. Suppose that $\{x^{k}\}$ has a cluster point $z\notin{\rm SOL}(F,C)$. Since SOL$(F,C)$ is closed by Proposition 2.7, and $\lim_{k\to\infty}\left\lVert x^{k+1}-x^{k}\right\rVert=0$ by Lemma 3.9(iii), we invoke Proposition 2.10 to obtain a subsequence $\{x^{j_{k}}\}$ of $\{x^{k}\}$ and a real number $\sigma>0$ such that
${\rm dist}\left(x^{j_{k}},{\rm SOL}(F,C)\right)>\sigma,$ (3.32)
and
${\rm dist}\left(x^{j_{k}+1},{\rm SOL}(F,C)\right)>{\rm dist}\left(x^{j_{k}},{\rm SOL}(F,C)\right).$ (3.33)
Take $\gamma_{k}$ as defined in (3.26). Note that for $\bar{x}\in{\rm SOL}(F,C)$, $\{\gamma_{k}(\bar{x})\}$ is bounded by Lemma 3.9(ii). Define $\gamma:{\rm SOL}(F,C)\to\mathds{R}$ as
$\gamma(x):=\liminf_{k\rightarrow\infty}\gamma_{j_{k}}(x).$ (3.34)
We claim that $\gamma(x)>0$ for all $x\in{\rm SOL}(F,C)$. Otherwise, By Proposition 2.8 $\{x^{j_{k}}\}$ has a cluster point in SOL$(F,C)$, in contradiction with (3.32). We claim now that $\gamma$ is continuous in SOL(F,C). Take $x,x^{\prime}\in{\rm SOL}(F,C)$. Note that
$\gamma_{j_{k}}(x^{k})=\langle u^{j_{k}},x^{j_{k}}-x\rangle=\langle u^{j_{k}},x^{j_{k}}-x^{\prime}\rangle+\langle u^{j_{k}},x^{\prime}-x\rangle\leq\gamma_{j_{k}}(x^{\prime})+\theta\left\lVert x-x^{\prime}\right\rVert.$
Thus, $\gamma(x)\leq\gamma(x^{\prime})+\theta\left\lVert x-x^{\prime}\right\rVert$, where $\theta$ is an upper bound of $\{\left\lVert u^{k}\right\rVert\}$. Reversing the role of $x,x^{\prime}$, we obtain $\left\lvert\gamma(x)-\gamma(x^{\prime})\right\rvert\leq\theta\left\lVert x-x^{\prime}\right\rVert$, establishing the claim.
Let $V$ be the set of cluster points of $\{x^{k}\}$. We have shown above that $V\cap{\rm SOL(F,C)}\neq\emptyset$. Since $\{x^{k}\}$ is bounded, $V$ is compact and so is $V\cap{\rm SOL}(F,C)$. It follows that $\gamma$ attains its minimum on $V\cap{\rm SOL}(F,C)$ at some $x^{*}$, so that $\gamma(x)\geq\gamma(x^{*})>0$ for all $x\in V\cap{\rm SOL}(F,C)$, using the claim above.
Take $\hat{k}$ such that
$\gamma_{j_{k}}(x)\geq\frac{\gamma(x^{*})}{2},$ (3.35)
and
$\beta_{j_{k}}<\frac{\gamma(x^{*})}{\theta},$ (3.36)
for all $k\geq\hat{k}$. Note that $\hat{k}$ exists because, for large enough $k$, (3.35) holds by virtue of (3.34), and (3.36) holds because $\lim_{k\to\infty}\beta_{k}=0$. In view of (3.27), we get, for all $x\in V\cap{\rm SOL}(F,C)$ and all $k\geq\hat{k}$,
$\begin{split}\left\lVert x^{j_{k}+1}-x\right\rVert^{2}&\leq\left\lVert x^{j_{k}}-x\right\rVert^{2}-\beta_{j_{k}}\left(2\frac{\gamma_{j_{k}}(x)}{\eta_{j_{k}}}-\beta_{j_{k}}\right)\\ &\leq\left\lVert x^{j_{k}}-x\right\rVert^{2}-\beta_{j_{k}}\left(2\frac{\gamma_{j_{k}}(x)}{\theta}-\beta_{j_{k}}\right)\\ &<\left\lVert x^{j_{k}}-x\right\rVert^{2},\end{split}$ (3.37)
using (3.35) in the second inequality and (3.36) in the third one. It follows that
${\rm dist}(x^{j_{k}+1},V\cap{\rm SOL}(F,C))\leq{\rm dist}(x^{j_{k}},V\cap{\rm SOL}(F,C)),$
for all $k\geq\hat{k}$, in contradiction with 3.33. The contradiction arises from assuming that $\{x^{k}\}$ has cluster points out of SOL$(F,C)$, and therefore all cluster points of $\{x^{k}\}$ solve VI$(F,C)$. ∎
We summarize the convergence sequence properties of Algorithm 1 in the following corollary.
Corollary 3.11.
If $T$ is paramonotone and coercive (in the sense of Assumption 3.5), and ${\rm SOL(F,C)}\neq\emptyset$, then any sequence $\{x^{k}\}$ generated by Algorithm 1 is bounded, $\lim_{k\to\infty}\left\lVert x^{k+1}-x^{k}\right\rVert=0$, and all cluster points of $\{x^{k}\}$ belong to SOL$(F,C)$. If VIP$(F,C)$ has a unique solution then the whole sequence $\{x^{k}\}$ converges to it.
Proof.
It follows from Lemma 3.9(i) and (iii), and Theorem 3.10. ∎
3.3 Algorithm 1 when the $C_{i}$’s are sublevel sets of convex functions
In this subsection, we consider the case when $C=\cap_{i=1}^{m}C_{i}$ and each set $C_{i}$ is of the form
$C_{i}=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\},$ (3.38)
where each $g_{i}:\mathds{R}^{n}\to\mathds{R}$ is convex. In principle, this form entails no loss of generality, because we can take $g_{i}(x)=\left\lVert P_{C_{i}}(x)-x\right\rVert^{2}$, but this case turns out to be computationally advantageous only when we have easily computable formulae for $g_{i}$ and its subdifferential. We remark that this is indeed the case in a very large number of applications. When the $C_{i}$’s are as in (3.38), we can use a very convenient operator $S^{i}$, which leads to explicit closed formulae for the iteration of Algorithm 1.
We define a separator operator for the set $C_{i}$ as
$S^{i}(x)=\{y\in\mathds{R}^{n}:g_{i}(x)+\langle u^{i},y-x\rangle\leq 0\},$ (3.39)
where $u^{i}$ is a subgradient of $g_{i}$ at $x$.
In connection with the sequence $\{x^{k}\}$ generated by Algorithm 1 we define $S^{i}_{k}:=S_{i}(x^{k})$, i.e.,
$S^{i}_{k}=\{y\in\mathds{R}^{n}:g_{i}(x^{k})+\langle u^{ik},y-x^{k}\rangle\leq 0\},$ (3.40)
where now $u^{ik}$ is a subgradient of $g_{i}$ at $x^{k}$. Finally, we define ${\mathbf{S}}_{k}\subset\mathds{R}^{nm}$ as
${\mathbf{S}}_{k}=S^{1}_{k}\times\dots\times S^{m}_{k}.$ (3.41)
This choice of the set ${\mathbf{S}}_{k}$ leads to the closed formulae presentation of Algorithm 1, namely:
$x^{k+1}=T_{\textbf{S}_{k},\textbf{D}}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)=x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})-\alpha_{k}w^{k},$ (3.42)
where
$\alpha_{k}=\frac{\sum_{i=1}^{m}\left\lVert v^{ik}\right\rVert^{2}}{m\left\lVert w^{k}\right\rVert^{2}},$ (3.43)
$w^{k}=\frac{1}{m}\sum_{i=1}^{m}v^{ik},$ (3.44)
$v^{ik}=\left[\frac{\max\left\{0,g_{i}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)\right\}}{\left\lVert u^{ik}\right\rVert}\right]u^{ik},$ (3.45)
with $u^{ik}\in\partial g_{i}\left(x^{k}-\frac{\beta_{k}}{\eta_{k}}F(x^{k})\right)$.
Remark 3.12.
We present some comments on the formulae (3.42)–(3.45).
  • i)
    The formula of $\alpha_{k}$ follows from the definition of the circumcenter, using Proposition 2.2 in (3.45); the computation of $\alpha_{k}$ is similar to the one presented in [11], in the definition of the PACA Algorithm for solving a CFP. In fact, the main difference between our closed formulae above and those in [11] is the presence of the term $(\beta/\eta_{k})F(x^{k})$ in (3.42), related to the operator $F$, which is specific of VIP$(F,C)$ and is absent in CFP.
  • ii)
    The closed formulae (3.42)-(3.45) were not used in the convergence analysis of Subsection 3.2, which is performed for general separating sets satisfying the conditions in Definition 1.2, but we note that they are essential for the effective implementation of the algorithm.
  • iii)
    These closed formulae hold only when the sets $C_{i}$ are of the form $C_{i}=\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\}$ for convex functions $g_{i}$. The iteration formula (3.3), which is the one used in our convergence analysis, covers the case of general sets $C_{i}$ and the use of other separating operators ${\textbf{S}}_{k}$, including even the exact projection case, for which $S_{i}^{k}=C_{i}$ for all $i,k$.
  • iv)
    Observe that $\partial g_{i}(x)\neq\emptyset$ for all $x\in\mathds{R}^{n}$, because we assume that $g_{i}$ is convex and dom$(g_{i})=\mathds{R}^{n}$ for all $i$.
In order to ensure that the convergence analysis in Subsection 3.2 holds for ${\mathbf{S}}_{k}$ as defined by (3.40), (3.41), we must prove that the separating operator $S^{i}$ defined by (3.39) satisfies indeed the conditions in Definition 1.2. We establish this in the next two propositions.
Proposition 3.13.
Let $f:\mathds{R}^{n}\to\mathds{R}$ be a convex function and $V\subset\mathds{R}^{n}$a nonempty, convex and compact set. Define $C=\{z\in\mathds{R}^{n}:f(z)\leq 0\}$. Given $x\in\mathds{R}^{n}$ and $v\in\partial f(x)$, define $\bar{S}(x,u):=\{z\in\mathds{R}^{n}|f(x)+\langle u,z-x\rangle\leq 0\}$. Then, there exists $\kappa>0$ such that
${\rm dist}(P_{\bar{S}(x,v)\cap X}(x),C)\leq\kappa\,{\rm dist}(x,C).$ (3.46)
for all $x\in V$ and all $u\in\partial f(x)$.
Proof.
See Lemma 4 in [32]. ∎
Proposition 3.14.
Given a convex function $g:\mathds{R}^{n}\to\mathds{R}\,\,(1\leq i\leq m)$ define $C:=\{x\in\mathds{R}^{n}:g(x)\leq 0\}$. Define $S(x)\subset\mathds{R}^{n}$ as
$S(x):=\{y\in\mathds{R}^{n}:g(x)+\langle u,y-x\rangle\leq 0\},$
where $u\in\mathds{R}^{n}$ is a subgradient of $g$ at $x$. Then $S$ is a separating set for $C$ in the sense of Definition 1.2.
Proof.
The fact that $S$ satisfies conditions (i)-(iii) in Definition 1.2 has been proved in Proposition 2.8 of [3]. We now deal with condition (iv). The result is an immediate consequence of Proposition 3.13. ∎
Proposition 3.14 establishes that ${\mathbf{S}}_{k}$, as defined in (3.41), fits the convergence analysis of Algorithm 1 carried out in Subsection 3.2.
We mention that an argument similar to the one in the proof of Proposition 3.14 establishes that the operator ${\mathbf{S}}:\mathds{R}^{nm}\to{\mathcal{P}}(\mathds{R}^{nm})$ given by ${\mathbf{S}}({\mathbf{x}})=(S^{1}(x^{1}),\dots,S^{m}(x^{m}))$, where ${\mathbf{x}}=(x^{1},\dots,x^{m})$ with $x^{i}\in\mathds{R}^{n}\,\,(1\leq i\leq m)$, and $S^{i}$ is defined as in (3.39), satisfies conditions (i)-(iv) in Definition 1.2. The proof for conditions (i)-(iii) can be found in [3]. However, this result is not needed for our convergence analysis.
4 An explicit circumcenter method for monotone variational inequality problem
In this section, we introduce an algorithm that utilizes approximate projections (projections onto halfspaces) instead of the exact projection. Different from the algorithm in Section 3, this algorithm only requires the assumption of monotonicity instead of paramonotonicity. First, we introduce the Explicit Circumcenter Method for monotone variational inequalities.
Algorithm 2 Explicit Circumcenter Method (ECM) for ${\rm VIP}(F,C)$ with monotone operators
Initialization:
Fix an exogenous constant $\theta>0$ and define
$x^{0}:=0~{}and~{}z^{0}\in\mathds{R}^{n}$
Choose a stepsize sequence $\{\beta_{k}\}$ satisfying:
$\sum_{k=0}^{\infty}\beta_{k}=\infty,$ (4.1)
and
$\sum_{k=0}^{\infty}\beta_{k}^{2}<\infty,$ (4.2)
Iterative Step:
Given $z^{k}$, if $g_{i}(z^{k})\leq 0$ for all $i\in\{1,\ldots,m\}$, then take $\tilde{y}^{k}:=z^{k}$. Else, perform the following inner loop, generating points $y^{k,0}$, $y^{k,1}$, $\ldots$. Take $y^{k,0}=z^{k}$, define
$y^{k,j+1}:=T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j}).$ (4.3)
Stop the inner loop when $j=j(k)$ , defined as
$j(k):=\min\left\{j\geq 0:\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}\leq\theta\beta_{k}\right\},$ (4.4)
where
$g(x)=\max_{1\leq i\leq m}g_{i}(x).$ (4.5)
Set
$\tilde{y}^{k}:=y^{k,j(k)}.$ (4.6)
Define $\eta_{k}:=\max\{1,\|T(\tilde{y}^{k})\|\}$. Take
$z^{k+1}:=T_{\mathbf{S}_{k},\mathbf{D}}\left(\tilde{y}^{k}-\frac{\beta_{k}}{\eta_{k}}T(\tilde{y}^{k})\right).$ (4.7)
If $z^{k+1}=y^{k}$, stop. Otherwise, define
$\sigma_{k}:=\sum_{j=0}^{k}\frac{\beta_{k}}{\eta_{k}}=\sigma_{k-1}+\frac{\beta_{k}}{\eta_{k}},$ (4.8)
$x^{k+1}:=\left(1-\frac{\beta_{k}}{\eta_{k}\sigma_{k}}\right)x^{k}+\frac{\beta_{k}}{\eta_{k}\sigma_{k}}\tilde{y}^{k}.$ (4.9)
We assume in this section that there exist a Slater point $w\in\mathds{R}^{n}$. This assumption is only needed in the inner loop (4.3) in order to check whether the current iterate in the loop is close enough to the feasible set.
4.1 Analysis of the direct method
For the convergence of our method, we assume that $T$ is monotone and continuous. Observe that $\partial g_{i}(x)\neq\emptyset$ for $x\in\mathds{R}^{n}$ for all $i\in\{1,\ldots,m\}$, because we assume that $g_{i}$ is convex and dom$(g_{i})=\mathds{R}^{n}$, for all $i\in\{1,\ldots,m\}$. First, we establish that Algorithm A is well-defined.
Lemma 4.1.
Take $C_{i}$, $\tilde{y}^{k}$, $z^{k}$ and $x^{k}$ defined by (2.3), (4.6), (4.7) and (4.9) respectively. Then
  • i)
    $g(y^{k,j})+\langle v^{k,j},y^{k,j+1}-y^{k,j}\rangle\leq 0$, where $g$ is defined in (4.5);
  • ii)
    If $z^{k+1}=\tilde{y}^{k}$, then $\tilde{y}^{k}\in{\rm SOL}(F,C)$.
  • iii)
    $j(k)$ is well defined.
Proof.
  • i)
    By the definition of $y^{k,j+1}$ and Proposition 2.3, we know that $y^{k+1}\in C_{i}$ for all $i\in\{1,\ldots,m\}$. According to the definition of $C_{i}$ and $g$, we know that (i) holds automatically.
  • ii)
    Suppose that $z^{k+1}=\tilde{y}^{k}$. Then, by Proposition 2.3, we get that $\mathbf{\tilde{y}}^{k}\in\mathbf{H}^{k}\subset\mathbf{S}^{k}\subset\mathds{R}^{nm}$. Therefore, $\tilde{y}^{k}\in S_{i}^{k}$ for all $i\in\{1,\ldots,m\}$. Given $i\in\{1,\ldots,m\}$, we have two cases, according to the definition of $S_{i}^{k}$ in (2.4). If $S_{i}^{k}=C_{i}$, then we have $z^{k+1}=\tilde{y}^{k}\in C_{j}$. Otherwise, we have $g_{j}(\tilde{y}^{k})\leq g_{j}(\tilde{y}^{k})+\langle v^{jk},z^{k+1}-\tilde{y}^{k}\rangle\leq 0$, i.e., $\tilde{y}^{k}\in C_{j}$. By the above argument above, we conclude that for any $i\in\{1,\ldots,m\}$, we have $\tilde{y}^{k}\in C_{i}$.
  • iii)
    Assume by contradiction that $\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}>\theta\beta_{k}$ for all $j$. Thus, we obtain an infinite sequence $\{y^{k,j}\}$ such that
    $\liminf_{j\rightarrow\infty}\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}\geq\theta\beta_{k}>0.$ (4.10)
    Taking into account the inner loop in $j$ given in (4.3) i.e., $y^{k,j+1}:=T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})$ for each fixed $k$, we obtain, for each $x\in C$,
    $\begin{split}\|y^{k,j+1}-x\|^{2}&=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})-T_{\mathbf{S}_{k},\mathbf{D}}(x)\|^{2}=\|P_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})-P_{\mathbf{S}_{k},\mathbf{D}}(x)\|\\ &\leq\|y^{k,j}-x\|^{2}-\|y^{k,j+1}-y^{k,j}\|^{2}\\ &\leq\|y^{k,j}-x\|^{2},\end{split}$ (4.11)
    where the first equality follows from Proposition 2.3 and the first inequality holds because of the firm nonexpansiveness of the projection operator. Thus, $\{y^{k,j}\}_{j\geq 0}$ is a quasi-Fejér convergent to $C$, and hence it is bounded by Proposition 2.15. It follows that
    $\tau:=\frac{1}{-g(w)}\sup_{0\leq j\leq\infty}\|y^{k,j}-w\|,$ (4.12)
    is finite, because $w$ is a Slater point. Also,
    $g(y^{k,j})>0~{}for~{}all~{}j.$ (4.13)
    Using (4.11), we get
    $\lim_{j\rightarrow\infty}\|y^{k,j+1}-y^{k,j}\|=\lim_{j\rightarrow\infty}\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j})-y^{k,j}\|=0.$ (4.14)
    By Lemma 4.1(i), we have
    $g(y^{k,j})\leq\langle v^{k,j},y^{k,j}-y^{k,j+1}\rangle\leq\|v^{k,j}\|\|y^{k,j}-y^{k,j+1}\|.$ (4.15)
    Since $\{y^{k,j}\}_{j\geq 0}$ is bounded, $g$ is convex, and $\partial g$ is bounded on bounded sets, we obtain that $\{v^{k,j}\}_{j\geq 0}$ is bounded. In view of (4.14) and (4.15),
    $\liminf_{j\rightarrow\infty}g(y^{k,j})\leq 0.$ (4.16)
    It follows from (4.13) and (4.16) that
    $\begin{split}\liminf_{j\rightarrow\infty}\frac{g(y^{k,j})\|y^{k,j}-w\|}{g(y^{k,j})-g(w)}&\leq\liminf_{j\rightarrow\infty}\frac{g(y^{k,j})\|y^{k,j}-w\|}{-g(w)}\\ &\leq\frac{1}{-g(w)}\sup_{0\leq j\leq\infty}\|y^{k,j}-w\|\liminf_{j\rightarrow\infty}g(y^{k,j})\\ &=\tau\liminf_{j\rightarrow\infty}g(y^{k,j})\leq 0,\end{split}$ (4.17)
    contradicting (4.10). Therefore, $j(k)$ is well defined.
Lemma 4.2.
If ${\rm SOL}(F,C)$ is nonempty, then $\{\tilde{y}^{k}\}$ and $\{z^{k}\}$ are quasi-Fejer convergent to ${\rm SOL}(F,C)$.
Proof.
Observe that $\eta_{k}\geq\|F(\tilde{y}^{k})\|$ and $\eta_{k}\geq 1$ for all $k$ by the definition of $\eta_{k}$. Then, for all $k$,
$\frac{1}{\eta_{k}}\leq 1$ (4.18)
and
$\frac{\|F(\tilde{y}^{k})\|}{\eta_{k}}\leq 1.$ (4.19)
Take $x^{*}\in{\rm SOL}(F,C)$. Then,
$\begin{split}\|\tilde{y}^{k}-x^{*}\|^{2}&=\|y^{k,j(k)}-x^{*}\|^{2}=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j(k-1)})-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\|^{2}\leq\|y^{k,j(k)-1}-x^{*}\|^{2}\\ &=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,j(k)-2})-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\|^{2}\leq\|y^{k,j(k)-2}-x^{*}\|^{2}\\ &\cdots\\ &\leq\|y^{k,1}-x^{*}\|^{2}=\|T_{\mathbf{S}_{k},\mathbf{D}}(y^{k,0})-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\|^{2}\\ &\leq\|y^{k,0}-x^{*}\|^{2}=\|z^{k}-x^{*}\|^{2}.\end{split}$ (4.20)
Let $\tilde{\theta}=1+\theta\|F(x^{*})\|\geq 1+\theta\frac{\|F(x^{*})\|}{\eta_{k}}$, by (4.18). Then
$\begin{split}\|\tilde{y}^{k+1}-x^{*}\|^{2}&\leq\|z^{k+1}-x^{*}\|^{2}=\left\|T_{\mathbf{S}_{k},\mathbf{D}}(\tilde{y}^{k}-\frac{\beta_{k}}{\eta_{k}}F(\tilde{y}^{k}))-T_{\mathbf{S}_{k},\mathbf{D}}(x^{*})\right\|^{2}\leq\|\tilde{y}^{k}-\frac{\beta_{k}}{\eta_{k}}F(\tilde{y}^{k})-x^{*}\|^{2}\\ &=\|\tilde{y}^{k}-x^{*}\|^{2}+\frac{\|F(\tilde{y}^{k})\|^{2}}{\eta_{k}^{2}}\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(\tilde{y}^{k}),\tilde{y}^{k}-x^{*}\rangle\\ &\leq\|\tilde{y}^{k}-x^{*}\|^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\langle F(x^{*}),\tilde{y}^{k}-x^{*}\rangle\\ &=\|\tilde{y}^{k}-x^{*}\|^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}(\langle F(x^{*}),\tilde{y}^{k}-P_{C}(\tilde{y}^{k})\rangle+\langle F(x^{*}),P_{C}(\tilde{y}^{k})-x^{*}\rangle)\\ &\leq\|\tilde{y}^{k}-x^{*}\|^{2}+\beta_{k}^{2}-2\frac{\beta_{k}}{\eta_{k}}\|F(x^{*})\|\|P_{C}(\tilde{y}^{k})-\tilde{y}^{k}\|\\ &\leq|\tilde{y}^{k}-x^{*}\|^{2}+\tilde{\theta}\beta_{k}^{2}\\ &\leq\|z^{k}-x^{*}\|^{2},\end{split}$ (4.21)
using (4.20) in the last inequality, the nonexpansiveness of $T_{\mathbf{S}_{k},\mathbf{D}}$ in the second one, the monotonicity of $F$ and (4.19) in the third one, $x^{*}\in{\rm SOL}(F,C)$ in the fourth one, Cauchy-Schwartz inequality in the fifth one, Lemma 2.18 and definition of $j(k)$ in the sixth one, and (4.20) in the last one.
In view of Definition 2.12, we conclude that $\{\tilde{y}^{k}\}$ and $\{z^{k}\}$ are quasi-Fejér convergent to ${\rm SOL}(F,C)$. ∎
Lemma 4.3.
Let $\{z^{k}\}$, $\{\tilde{y}^{k}\}$ and $\{x^{k}\}$ be the sequences generated by Algorithm 2. Assume that ${\rm SOL}(F,C)$ is nonempty. Then
  • i)
    $x^{k+1}=\frac{1}{\sigma_{k}}\sum_{k=1}^{k}\frac{\beta_{j}}{\eta_{j}}\tilde{y}^{j}$;
  • ii)
    $\{\tilde{y}^{k}\}$, $\{x^{k}\}$ and $\{F(\tilde{y}^{k})\}$ are bounded;
  • iii)
    $\lim_{k\rightarrow\infty}dist(x^{k},C)=0$;
  • iv)
    all the cluster points of $\{x^{k}\}$ belong to $C$.
Proof.
  • i)
    Note that (4.8) and (4.9) can be rewritten as
    $\begin{split}\sigma_{k}x^{k+1}=\left(\sigma_{k}-\frac{\beta_{k}}{\eta_{k}}\right)x^{k}+\frac{\beta_{k}}{\eta_{k}}\tilde{y}^{k}&=\sigma_{k-1}x^{k}+\frac{\beta_{k}}{\eta_{k}}\tilde{y}^{k}\\ &=\sigma_{k-2}x^{k-1}+\frac{\beta_{k-1}}{\eta_{k-1}}\tilde{y}^{k-1}+\frac{\beta_{k}}{\eta_{k}}\tilde{y}^{k}\\ &\cdots\\ &=\sigma_{0}x^{1}+\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\tilde{y}^{j}\\ &=\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\tilde{y}^{j},\end{split}$ (4.22)
    using the fact that $\sigma_{0}=0$, which follows from (4.8). We obtain the result after dividing both sides by $\sigma_{k}$.
  • ii)
    For $\{\tilde{y}^{k}\}$ we use Lemma 4.2 and Proposition 2.15(i). For $\{F(\tilde{y}^{k})\}$, we use the boundedness of $\{\tilde{y}^{k}\}$ and the fact that $F$ is monotone and continuous. For $\{x^{k}\}$, use the boundedness of $\{\tilde{y}^{k}\}$ and (i).
  • iii)
    By Lemma 2.18 and (4.4)-(4.6), we have
    $dist(\tilde{y}^{k},C)\leq\theta\beta_{k}.$ (4.23)
    Define
    $\tilde{x}^{k+1}:=\frac{1}{\sigma_{k}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}P_{C}(\tilde{y}^{k}).$ (4.24)
    Since $\frac{1}{\sigma_{k}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}=1$ by (4.8), we get from the convexity of $C$ that $\tilde{x}^{k+1}\in C$. Let
    $\tilde{\beta}:=\sum_{j=0}^{\infty}\beta_{j}^{2}.$ (4.25)
    Note that $\tilde{\beta}$ is finite by (4.1). Then
    $\begin{split}dist(x^{k+1},C)&\leq\|x^{k+1}-\tilde{x}^{k+1}\|=\left\|\frac{1}{\sigma_{k}}\left(\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}(\tilde{y}^{j}-P_{C}(\tilde{y}^{j}))\right)\right\|\\ &\leq\frac{1}{\sigma_{k}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\|\tilde{y}^{j}-P_{C}(\tilde{y}^{j})\|\\ &\leq\frac{\theta}{\sigma_{k}}\sum_{j=0}^{k}\frac{\beta_{j}^{2}}{\eta_{j}}\leq\frac{\theta}{\sigma_{k}}\sum_{j=0}^{k}\beta_{j}^{2}\leq\theta\frac{\tilde{\beta}}{\sigma_{k}},\end{split}$ (4.26)
[…]
    Take $\gamma>1$ such that $\|F(\tilde{y}^{k})\|\leq\gamma$ for all $k$. Note that existence of $\gamma$ follows from (i). Then,
    $\lim_{k\rightarrow\infty}\sigma_{k}=\lim_{k\rightarrow\infty}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\geq\lim_{k\rightarrow\infty}\frac{1}{\gamma}\sum_{j=0}^{k}\beta_{j}=\infty,$ (4.27)
    using the fact that $\eta_{j}=max\{1,\|F(\tilde{y})\|\}\leq max\{1,\gamma\}\leq\gamma$ for all $j$ in the inequality, and (4.1) in the first equality. Thus, taking limits in (4.26), we get
    $\lim_{k\rightarrow\infty}dist(x^{k},C)=0,$
    establishing (iii). Item (iv) follows immediately from (iii).
Theorem 4.4.
If ${\rm SOL}(F,C)$ is nonempty then all the cluster points of $\{x^{k}\}$ generated by Algorithm 2 are solutions of ${\rm VIP}(F,C)$.
Proof.
For any $x\in C$ we have
$\begin{split}\|z^{j+1}-x\|&=\left\|T_{\mathbf{S}_{k},\mathbf{D}}\left(\tilde{y}^{j}-\frac{\beta_{j}}{\eta_{j}}F(\tilde{y}^{j})\right)-T_{\mathbf{S}_{k},\mathbf{D}}(x)\right\|\leq\left\|\left(\tilde{y}^{j}-\frac{\beta_{j}}{\eta_{j}}F(\tilde{y}^{j})\right)-x\right\|\\ &=\|\tilde{y}^{j}-x\|^{2}+\frac{\|F(\tilde{y}^{j})\|}{\eta_{j}^{2}}\beta_{j}^{2}-2\frac{\beta_{j}}{\eta_{j}}\langle F(\tilde{y}^{j}),\tilde{y}^{j}-x\rangle\\ &\leq\|\tilde{y}^{j}-x\|^{2}+\beta_{j}^{2}+2\frac{\beta_{j}}{\eta_{j}}\langle F(x),x-\tilde{y}^{j}\rangle,\end{split}$ (4.28)
using nonexpansiveness of $T_{\mathbf{S}_{k},\mathbf{D}}$ in the first inequality, and monotonicity of $F$ and (4.19) in the last inequality. Summing (4.28) from 00 to $k-1$ and dividing by $\sigma_{k-1}$, we obtain, in view of Lemma 4.3(ii),
$\frac{(\|z^{k}-x\|^{2}-\|z^{0}-x\|^{2})}{\sigma_{k-1}}\leq\frac{\sum_{j=0}^{k-1}\beta_{j}^{2}}{\sigma_{k-1}}+\langle F(x),x-x^{k}\rangle.$ (4.29)
Let $\hat{x}$ be a cluster point of $\{x^{k}\}$. Existence of $\hat{x}$ is guaranteed by Lemma 4.3(i). Note that $\hat{x}\in C$ by Lemma 4.3(iv).
By (4.1), (4.2), (4.27) and boundedness of $\{z^{k}\}$, taking limits in (4.29) with $k\rightarrow\infty$, we obtain that
$\langle F(x),x-\hat{x}\rangle\geq 0,$
for all $x\in C$. Using Proposition 2.14, we get that $\hat{x}\in{\rm SOL}(F,C)$. Therefore, all the cluster points of $\{x^{k}\}$ solve ${\rm VIP}(F,C)$. ∎
Theorem 4.5.
Define $\bar{x}=\lim_{k\rightarrow\infty}P_{{\rm SOL}(F,C)}(\tilde{y}^{k})$. Then, either ${\rm SOL}(F,C)\neq\emptyset$ and $\{x^{k}\}$ converges to $\bar{x}$, or ${\rm SOL}(F,C)=\emptyset$ and $\lim_{k\rightarrow\infty}\|x^{k}\|=\infty$.
Proof.
Assume that ${\rm SOL}(F,C)\neq\emptyset$ and define $u^{k}=P_{{\rm SOL}(F,C)}(\tilde{y}^{k})$. Note that $u^{k}$ exists because ${\rm SOL}(F,C)$ is nonempty, closed, and convex by Proposition 2.13. By Lemma 4.2, $\{\tilde{y}^{k}\}$ is quasi-Fejér convergent to ${\rm SOL}(F,C)$ since ${\rm SOL}(F,C)\subset C$. Therefore, it follows from Proposition 2.16(ii) that $\{P_{{\rm SOL}(F,C)}(\tilde{y}^{k})\}$ is convergent. Let
$\bar{x}=\lim_{k\rightarrow\infty}P_{{\rm SOL}(F,C)}(\tilde{y}^{k})=\lim_{k\rightarrow\infty}u^{k}.$ (4.30)
By Lemma 4.2 and Theorem 4.4, $\{x^{k}\}$ is bounded and each of its cluster points belongs to ${\rm SOL}(F,C)$. Suppose that $\hat{x}\in{\rm SOL}(F,C)$ is a cluster point of $\{x^{k}\}$ and $\{x^{i_{k}}\}\rightarrow\hat{x}$. It suffices to show that $\bar{x}=\hat{x}$.
By Proposition 2.1(i) we have that $\langle\hat{x}-u^{j},\tilde{y}^{j}-u^{j}\rangle\leq 0$ for all $j$. Let $\epsilon=\sup_{0\leq j\leq\infty}\|\tilde{y}^{j}-u^{j}\|$. By Lemma 4.3(i), $\epsilon<\infty$. Then,
$\langle\hat{x}-\bar{x},\tilde{y}^{j}-u^{j}\rangle\leq\langle u^{j}-\bar{x},\tilde{y}^{j}-u^{j}\rangle\leq\epsilon\|u^{j}-\bar{x}\|,$ (4.31)
for all $j$. Multiplying (4.31) by $\frac{\beta_{j}}{\eta_{j}\sigma_{k-1}}$ and summing from 00 to $k-1$, we get from Lemma 4.3(ii)
$\left\langle\hat{x}-\bar{x},x^{k}-\frac{1}{\sigma_{k-1}}\sum_{j=0}^{k-1}\frac{\beta_{j}}{\eta_{j}}u^{j}\right\rangle\leq\frac{\epsilon}{\sigma_{k-1}}\sum_{j=0}^{k-1}\frac{\beta_{j}}{\eta_{j}}\|u^{j}-\bar{x}\|.$ (4.32)
Define $\zeta_{k,j}:=\frac{1}{\sigma_{k}}\frac{\beta_{j}}{\eta_{j}}(k\geq 0,~{}0\leq j\leq k)$. In view of (4.27), $\lim_{k\rightarrow\infty}\zeta_{k,j}=0$ for all $j$. By (4.30),
$\sum_{j=0}^{k}\zeta_{k,j}=1,$ (4.33)
for all $k$. Then, using (4.30) and Proposition 2.17 with
$w^{k}=\sum_{j=0}^{k}\zeta_{k,j}u^{j}=\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}u^{j},$ (4.34)
we obtain that
$\lim_{k\rightarrow\infty}\frac{1}{\sigma_{k-1}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}u^{j}=\bar{x},$ (4.35)
and hence
$\lim_{k\rightarrow\infty}\frac{1}{\sigma_{k-1}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}\|u^{j}-\bar{x}\|=0,$ (4.36)
using the fact that $\frac{1}{\sigma_{k-1}}\sum_{j=0}^{k}\frac{\beta_{j}}{\eta_{j}}=1$.
From (4.35) and (4.36), since $\lim_{k\rightarrow\infty}x^{i_{j}}=\bar{x}$, taking limits in (4.34) with $k\rightarrow\infty$ along the subsequence with subindices $\{i_{k}\}$, we get that
$\langle\hat{x}-\bar{x},\hat{x}-\bar{x}\rangle\leq 0,$
implying that $\hat{x}=\bar{x}$.
If ${\rm SOL}(F,C)$ is empty then by Theorem 4.4 no subsequence of $\{x^{k}\}$ can be bounded, and hence $\lim_{k\rightarrow\infty}\|x^{k}\|=\infty$. ∎
We remark that the sequence $\{x^{k}\}$ is a weighted average of the $k$ first elements of the sequence $\{\tilde{y}^{k}\}$. It is easy to check that if $\{\tilde{y}^{k}\}$ converges then so does $\{x^{k}\}$, but the converse statement does not necessarily hold. Indeed, the basic sequence in this algorithms is $\{\tilde{y}^{k}\}$, and the sequence $\{x^{k}\}$ was introduced only because we were not able to prove directly the convergence of $\{\tilde{y}^{k}\}$ ($\{x^{k}\}$ can be considered as an ergodic sequence related to $\{\tilde{y}^{k}\}$). In all our numerical experiments, the sequence $\{\tilde{y}^{k}\}$ did indeed converge, and obviously, faster than $\{x^{k}\}$, whose convergence is slowed down by the effect of the initial elements of $\{\tilde{y}^{k}\}$ in the computation of $x^{k}$. Hence, in Section 5 we report the results related to $\{\tilde{y}^{k}\}$ instead of $\{x^{k}\}$.
5 Numerical experiments
In this section, we present numerical experiments which evaluate the performance of our proposed Algorithms 1 and 2. We compare them with Algorithm 2 in [17], Algorithm 1 in [18], and the classical Extragradient method [45] for solving the variational inequality problem ${\rm VIP}(F,C)$. For clarity, we refer to these algorithms as Alg1, Alg2, Alg3, Alg4, and Extra, respectively. We recall that Alg1 is a circumcenter-based method intended to accelerate Alg3, and Alg2 acts in a similar way with respect to Alg4. Algorithms Alg1–Alg4 employ approximate projections onto convex sets, while Extra uses exact projections.
We consider the feasible set $C$ defined as the intersections of ellipsoids:
$C=\bigcap_{i=1}^{m}C_{i}.\quad C_{i}=\left\{x\in\mathds{R}^{n}:g_{i}(x)\leq 0\right\}=\left\{x\in\mathds{R}^{n}:x^{\top}A_{i}x+2(b^{i})^{\top}x-\alpha\leq 0\right\},$ (5.1)
where each $A_{i}\in\mathds{R}^{n\times n}$ is a positive definite matrix, $b^{i}\in\mathds{R}^{n}$, and $\alpha\in\mathds{R}$. These ellipsoids are generated following the approach in [10, Section 5], thus ensuring the existence of a Slater point, required for the inner loop of both Alg2 and Alg4.
For computing the circumcenter in (1.23), we use the gradient of $g_{i}$, which is given by $\nabla g_{i}(x)=2A_{i}x+b^{i}$. However, computing the exact projection onto $\bigcap_{i=1}^{m}C_{i}$ required in the Extragradient steps (1.6) and (1.7) is nontrivial. We employ Dykstra’s algorithm [29] and ADMM [41, Algorithm 6] for computing the projection onto each $C_{i}$, which turn out to work well for this purpose.
Since Alg2 [17] and Alg4 [18] require halfspaces separating $C$ from the current iterate, we define the function
$\hat{g}(x):=\max_{i\in\{1,\ldots,m\}}g_{i}(x),$ (5.2)
and construct the halfspace $\hat{H}(x)$, defined as
$\hat{H}(x):=\left\{z\in\mathds{R}^{n}:\hat{g}(z)\geq\hat{g}(x)+\langle u,z-x\rangle\right\},$ (5.3)
with $u\in\partial\hat{g}(x)$. It is well known that $u$ can be taken as $\nabla g_{k}(x)$, where $k$ is any index which realizes the maximum in (5.2). We mention that $\hat{H}(x)$ separates the point $x$ from the feasible set $C$.
We perform two kinds of experiments:
  • Approximate Projection Comparison: We compare Alg1–Alg4 in terms of CPU time and iteration counts, for highlighting the acceleration effect of the circumcenter approach in Alg1 and Alg2, with respect to the other two algorithms.
  • Exact vs. Approximate Projection: We compare Alg1, Alg2, and Extra in terms of CPU time and final error, in order to investigate the practical impact of using exact versus approximate projections.
The computational experiments were performed on an Intel Core i7-12700H 2.30GHz with 32 GB of RAM running Windows 10 Professional version using Matlab 2015b, and are fully available at
https://github.com/Yanziszj/CRM-VIP. The following conditions were used:
  1. 1.
    The maximum of iterations for all the algorithms is 30000.
  2. 2.
    The stopping criteria for Alg1 and Alg3 are
    $\|x^{k+1}-x^{k}\|\leq\epsilon~{}{\rm or}~{}k>30000.$ (5.4)
    The stopping criteria for Alg2 and Alg4 are
    $\|z^{k+1}-y^{k}\|\leq\epsilon~{}{\rm or}~{}k>30000.$ (5.5)
    The stopping criteria for Extra are
    $\|x^{k}-y^{k}\|\leq\epsilon~{}{\rm or}~{}k>30000,$ (5.6)
    where $\epsilon=10^{-6}$ and $k$ is the iteration index.
  3. 3.
    The parameters in ADMM are the same as in [41, Algorithm 6]. The stopping criterion of Dykstra’s Algorithm is the same as in [19, Sect. 4]
  4. 4.
    For determinig whether algorithms using approximate projection properly approach the solution set of ${\rm VIP}(F,C)$, we compute the error at iteration $k$ of each algorithm in the second part of our experiments by the following rule:
    ${\rm error}=\|x^{k}-P_{C}(x^{k}-\alpha F(x^{k}))\|,$ (5.7)
    with $\alpha>0$. We choose as $\alpha=0.1$. It is well known that if $x\in{\rm SOL}(F,C)$, then $x$ is also a fixed point of $\left[P_{C}\circ(I-\alpha F)\right](\cdot)$ for any positive $\alpha$.
  5. 5.
    For each pair $(n,m)$, for $n\in\{5,10,20\}$ and $m\in\{2,5,10\}$ we generate 20 examples and run the test for each example 10 times.
  6. 6.
    We choose $\beta_{k}=\frac{1}{k}$ for Alg1-Alg4 according to (3.4) and (3.5) in Algorithm 1, (4.1) and (4.2) in Algorithm and similar assumptions in [17] and [18]. We take a constant stepsize $\alpha=0.05$ for Extra.
5.1 Examples of operators in ${\rm VIP}(F,C)$
In order to test Alg1 and Ag2 for ${\rm VIP}(F,C)$ with paramonotone and monotone operators, we consider the following three families of operators.
Example 1: First, we test examples with operators $F$ which are gradients of convex functions. In these cases, ${\rm VIP}(F,C)$ is equivalent to the problem of minimizing the convex function with gradient $F$ over $C$. Define
$F(x)=Ax+G(x)+c$ (5.8)
where $G(x)_{i}=b_{i}x_{i}^{3}$ with $b,c\in\mathds{R}^{n},b\geq 0,c\neq 0$. We generate randomly $A\in\mathds{R}^{n\times n}$ as a symmetric, positive semidefinite matrix. Note that $F(x)$ is the gradient of the convex function $f(x)=\frac{1}{2}x^{t}Ax+cx+\sum_{i=1}^{n}b_{i}x_{i}^{4}$.
Example 2: It is known that the gradient of a convex function is a paramonotone operator [23] but the inverse is not true. We consider now operators which are not gradients of convex functions. These operators ae defined as:
$F(x)=Ax+c,$ (5.9)
where the matrix $A\subset\mathds{R}^{n\times n}$ is defined as
$A=\left[\begin{array}[]{cc}A_{1}&0\\ 0&A_{2}\end{array}\right].$ (5.10)
with a symmetric positive semidefinite matrix $A_{1}\in\mathds{R}^{n_{1}\times n_{1}}$ and a nonsymmetric positive semidefinite matrix $A_{2}\in\mathds{R}^{n_{2}\times n_{2}}$. $F$ is not a gradient, because $A$ is the Jacobian matrix of $F$ and Jacobian matrices of gradients are symmetric.
Example 3: Finally, we consider cases with monotone operators which are not paramonotone. $F$ is still defined by (5.9), and $A_{1}\in\mathds{R}^{n_{1}\times n_{1}}$ is also a symmetric positive semidefinite matrix, but $A_{2}\in\mathds{R}^{n_{2}\times n_{2}}$ in (5.10) is skew-symmetric, which means that
$A_{2}^{t}=-A_{2}$ (5.11)
Hence, ${\rm rank}(A+A^{t})<{\rm rank}(A)$. Note that $A$ is the Jacobian matrix of $F$. It is known that if $J(x)$ is the Jacobian matrix of a paramonotone differentiable operator, then ${\rm rank}(J+J^{t})={\rm rank}(J)$, so that the operator $F$ is not paramonotone in this case. Observe that $F$ is monotone, since operators with skew-symmetric Jacobian matrices are trivially monotone.
For each example, we list our results in two tables corresponding to both kinds of comparisons. Finally, we summary the result in all experiments in Figure 1.
Table 1: Median of CPU time and iteration number of Example 1
$n$ $m$ Alg1 Alg2 Alg3 Alg4
CPU(s) Iter CPU(s) Iter CPU(s) Iter CPU(s) Iter
5 2 0.031 908.5 0.141 1670 0.046 1045.5 0.156 1670.5
5 5 0.062 787 1.328 8499.5 1.164 15507 2.921 4320
5 10 0.132 860 8.414 30001 4.376 30001 14.078 10179
10 2 0.031 985 0.289 2991 0.062 1223 0.304 2991.5
10 5 0.063 880 3.250 19006 1.351 16278 1.734 4720
10 10 0.203 1336 7.257 21630 4.421 30001 44.445 26414
20 2 0.070 1519 0.304 2686.5 0.070 1442.5 0.320 2686.5
20 5 0.258 2995 5.085 27223 2.593 30001 20.898 30001
20 10 0.195 1190 3.398 9108.5 2.914 19215 43.087 26634
Table 2: Median of CPU time and error for Example 1
$n$ $m$ Alg1 Alg2 Extra
CPU (s) Error CPU (s) Error CPU (s) Error
5 2 0.031 0.01196 0.539 0.01195 0.523 0.00735
5 5 0.062 0.01026 1.500 0.01003 4.070 0.00724
5 10 0.125 0.00847 8.992 0.00802 23.796 0.00558
10 2 0.031 0.12329 0.187 0.01726 0.671 0.01127
10 5 0.156 0.01239 2.968 0.01244 29.843 0.00804
10 10 0.382 0.01239 9.031 0.01073 82.789 0.00788
20 2 0.054 0.01557 0.398 0.01507 3.085 0.01017
20 5 0.226 0.03118 1.835 0.03116 30.781 0.01906
20 10 0.421 0.03126 9.375 0.02990 171.515 0.02163
Table 3: Median of CPU time and iterations for Example 2
$n$ $m$ Alg1 Alg2 Alg3 Alg4
CPU(s) Iter CPU(s) Iter CPU(s) Iter CPU(s) Iter
5 2 0.031 738.5 0.195 2139.5 0.031 739 0.211 2139.5
5 5 0.062 857 1.906 10824 0.585 6818 7.859 9482
5 10 0.132 878.5 2.406 7737.5 4.429 30001 22.164 12450
10 2 0.031 782.5 0.382 93745 0.032 798.5 1.921 16543
10 5 0.101 1064 1.304 6597 0.250 2367.5 29.218 30001
10 10 0.187 1073 11.101 30001 5.492 30001 67.445 30001
20 2 0.109 3047 0.453 4502 0.125 3146.5 0.546 5093.5
20 5 0.445 2101 4.195 13305 1.945 18460 6.641 13415
20 10 0.437 2673 5.007 7739 53.609 24017 14.562 11119
Table 4: Median of CPU time and error for Example 2
$n$ $m$ Alg1 Alg2 Extra
CPU (s) Error CPU (s) Error CPU (s) Error
5 2 0.031 0.01298 0.321 0.01254 0.617 0.00804
5 5 0.109 0.00982 3.179 0.00962 11.226 0.00602
5 10 0.226 0.01178 9.242 0.01130 44.585 0.01091
10 2 0.148 0.01301 1.273 0.01307 4.968 0.00882
10 5 0.218 0.01431 3.882 0.01431 38.071 0.00932
10 10 0.414 0.01245 9.304 0.01240 122.117 0.00772
20 2 0.093 0.01592 0.375 0.01606 4.821 0.01106
20 5 0.375 0.02437 2.242 0.01948 88.812 0.01356
20 10 0.321 0.01495 8.234 0.01490 125.039 0.01082
Table 5: Median of CPU time and iterations for Example 3
$n$ $m$ Alg1 Alg2 Alg3 Alg4
CPU(s) Iter CPU(s) Iter CPU(s) Iter CPU(s) Iter
5 2 0.031 843.5 0.234 2648.5 0.039 843.5 0.250 2647.5
5 5 0.071 909 1.641 9876.5 2.046 25738 7.734 154749
5 10 0.156 1036 8.578 30001 4.289 30001 50.906 30001
10 2 0.046 1216 0.289 3199.5 0.054 1218 0.304 3199.5
10 5 0.062 762.5 1.195 7284.5 1.210 15399 4.109 5483.5
10 10 0.321 2189.5 8.601 30001 4.289 30001 50.906 30001
20 2 0.046 1394 0.273 2988 0.062 1434 0.281 2998
20 5 0.289 3672.5 1.632 8685 1.757 22185 18.203 21868
20 10 0.406 1577.5 14.125 30001 6.078 30001 70.023 30001
Table 6: Median of CPU time and error for Example 3
$n$ $m$ Alg1 Alg2 Extra
CPU (s) Error CPU (s) Error CPU (s) Error
5 2 0.031 0.01744 0.187 0.01740 0.328 0.01061
5 5 0.148 0.01248 4.234 0.01267 13.632 0.00749
5 10 0.141 0.00920 5.851 0.01658 21.054 0.00608
10 2 0.031 0.02538 0.211 0.02557 1.429 0.01484
10 5 0.078 0.02683 0.992 0.02684 13.828 0.01674
10 10 0.382 0.01663 7.820 0.01658 66.976 0.01062
20 2 0.141 0.15133 0.437 0.01677 6.890 0.01086
20 5 0.679 0.02091 3.031 0.02150 57.523 0.01360
20 10 1.085 0.02167 10.054 0.01919 207.695 0.01195
(a) CPU Time (Alg1-Alg4)
(b) Iteration number (Alg1-Alg4)
(c) CPU Time (Alg1-Alg2 and Extra)
(d) Errors (Alg1-Alg2 and Extra)
Figure 1: First row are comparison among Alg1-Alg4 and the second line are among Alg1-Alg2 and Extra.
5.2 Conclusion
Based on the results of the three types of numerical experiments, we can draw the following observations:
  1. 1.
    From Table 1 and Table 3, we observe that Alg1 and Alg3 perform similarly to Alg2 and Alg4 in terms of CPU time and number of iterations when $m=2$. However, as the number of ellipsoids increases ($m=5,10$), Alg1 and Alg3 significantly outperform Alg2 and Alg4, especially in the case when $m=10$, no matter the dimension $n$. This fact establishes unequivocally the superiority of the circumcenter methods with respect to the older algorithms when the number of ellipsoids increases.
  2. 2.
    From Table 2 and Table 4, the Extragradient method (Extra) demonstrates higher accuracy than both Alg1 and Alg2, which is not surprising, since Extra uses exact projections. Alg2, which includes an inner loop that approximates the exact projection, achieves better accuracy than Alg1. On the other hand, Alg1 achieves significantly faster computation times than Alg2, especially when the number of ellipsoids increases, while maintaining acceptable accuracy. Again, this faster performance confirms the superiority of circumcenter methods.
  3. 3.
    In Example 3, where the operator $F$ is monotone but no longer paramonotone, Alg3 fails to terminate within 30,000 iterations in some cases (see Table 6), consistent with the theoretical requirement of paramonotonicity for its convergence (as indicated in Table 5). In contrast, Alg1, which incorporates the circumcenter-acceleration technique, converges reliably and efficiently, with a reasonable number of iterations. While its accuracy is slightly lower compared to Extra, it is acceptable, and again, it performs much faster than Extra.
  4. 4.
    We also summarize our numerical findings in Fig 1, by means of the so-called performance profiles from [28]. Performance profiles allow one to benchmark different methods on a set of problems with respect to a performance measure (in our case, CPU time and total number of projections). The vertical axis indicates the percentage of problems solved, while the horizontal axis indicates the corresponding factor of the performance index used by the best solver.
In summary, Alg1 and Alg3 demonstrate superior performance over Alg2 and Alg4 in terms of CPU time and iteration count. Although Alg1 is slightly less accurate than Extra, due to the use of approximate projections, it offers a compelling balance between efficiency and accuracy, especially in problems with high-dimensional feasible sets.
6 Concluding remarks
In this work, we propose two distinct circumcenter-type projection methods for solving variational inequality problems involving paramonotone and monotone operators, respectively. Both algorithms enjoy global convergence and ergodic convergence guarantees. Numerical experiments demonstrate that our methods converge significantly faster than existing algorithms, especially when the intersection $\bigcap_{i=1}^{m}C_{i}$ is more complex, while maintaining reasonable accuracy.
Future research may focus on extending these methods to broader classes of operators, such as quasi-monotone, or stochastic settings, and investigating their applications in convex optimization or generalizing them in the equilibrium problem. Additionally, further theoretical analysis on convergence rates, and improving the ergodic convergence to global convergence of Algorithm 2 remain as important issues to be addressed.